### v3

# Extensions of $\mathcal{KL}$ and Lyapunov Functions for Discrete-time Dynamical System Peaks Analysis

Abstract
In this paper, we extend two classes of functions classically involved in asymptotic stability analyses for studying a maximization problem on the reachable values of a discrete-time dynamical system. This particular maximization problem is called a peaks computation problem. The problem is to find a couple composed of an initial state and a time which maximizes a given function over states. The paper focuses on the time component of the optimal solution which is an integer as the time is discrete. We develop a method to provide an upper bound of this integer from a formula which requires a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence. A first result proves that the formula provides, in theory, the optimal integer. However, in practice, the computation cannot be so precise. Then, we develop two alternative methods. The first is based on discontinuous and non strictly increasing/decreasing $\mathcal{KL}$-like functions named $\mathcal{KL}_{\rm gen}$ functions. We prove that the existence of a $\mathcal{KL}_{\rm gen}$ upper bound is equivalent to the existence of a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence. The construction of the strictly increasing continuous function from a $\mathcal{KL}_{\rm gen}$ function needs an extension of the famous Sontag’s lemma. Finally, we construct a new type of Lyapunov functions well designed for our peaks computation problem. These functions are called Opt-Lyapunov functions. Opt-Lyapunov functions are well designed since we establish an equivalence theorem between the existence of an Opt-Lyapunov function and of a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence. The construction of a Opt-Lyapunov function from a pair of a strictly increasing and continuous function on $[0,1]$ and a convergent geometric sequence is insipred by the Yoshizawa construction of Lyapunov functions.
Keywords: Discrete-time, Peaks Computation Problem, Supremum of real sequences, Lyapunov functions, $\mathcal{KL}$ functions.
1 Introduction
In this paper, we study the problem of peaks computation for discrete-time dynamical systems in finite dimensional state spaces. A peaks computation problem can be viewed as a particular maximization problem for which the constraints set is the reachable values set of a discrete-time dynamical system. The objective function of this optimization problem is a function over the states. This particular maximization problem can model various situations in engineering, economics, physics, biology… For example, in population dynamics, for a predator-prey system, one may be interested in the maximal size of each species even if the system asymptotically oscillates between several equilibra. This problem boils down to maximize separetely each coordinate of the state variable over the system modeling the population dynamics.
Besides its importance to analyze critical concrete situations, the literature about peaks analysis problem in the community of dynamical systems analysis is pretty thin. The peaks computation problem appears in a discrete-time setting for linear systems and where the objective function is the Euclidean norm [APS18]. The problem also appears for the analysis of stable linear systems with general quadratic objective functions [Adj21, Adj23]. A slighty different version of the problem is also encountered in [AG24]. The difference comes from the fact that the state variable is, moreover, constrained to stay in a some given set. The problem is studied in a continuous-time setting and in, algebraically, a polynomial environment in [MHS20]. Note that in [MHS20], the authors address the problem of verifying safety properties (see e.g. [BYG17] for verification problems on discrete-time systems) for dynamical systems and formulate the safety verification problem as a peaks computation problem. Safety verification problems aim to prove that the trajectories of the system are contained in some safe set. A safety verification problem can be translated into a peaks computation problem when the safe set is a sublevel of a sufficiently regular function. In this case, the objective function of the peaks computation problem is the function associated with the sublevel set. The safety analysis from the peaks computation point-of-view is also proposed for switched system [Adj17] and for program analysis in [AGM15]. For safety analysis of dynamical systems, one can think that a reachability analysis (see e.g. [RKML06] and reference therein) equivalent to a feasibility analysis in our context is sufficient to prove some properties on the dynamical system. However, maximality in our context is a synonym for criticality. The optimal value in this situation can be viewed as the value which penalizes the most the system with respect to some criteria and defines the extreme condition where the system could enter into.
In [MHS20] the authors are interested in computing an upper bound of the peaks optimal value for the continuous case. The techniques used is based on Sums-Of-Squares (SOS) and Linear Matrix Inequalities (LMI) formulations (the interested reader can consult [Las15] for those subjects). The same goal motivates the authors of [APS18] and, as they limit their study to linear system and norm objective functions, they directly use semidefinite programming (see [BTN01] for further details on the subject). The reference [AG24] cannot be fairly compared due to the presence of the additional constraint. Moreover, again, the studied systems are linear or piecewise linear and the objective functions are linear. The goal of the current paper is to provide the theoretical bases to solve exactly a peaks computation problem in a general way. The developments made in this paper can be viewed as a theoretical generalization of works done in [Adj21, Adj23]
In this paper, we propose to study peaks computation problems from classical tools of dynamical systems stability theory (e.g. [Ela05, Chap. 4] for discrete-time). This approach comes from a simple observation: the existence of peaks depends on the asymptotic behaviour of the system. The importance of stability analysis is time-tested and a stability analysis is a prerequisite to the implementation of numerical methods. Whereas general stability properties can be difficult to obtain, asymptotic stability properties can be easily presented and understood from $\mathcal{KL}$ bounds (e.g. [ATP10] or [NT04]) i.e. the norms of the states are globally bounded from above by a $\mathcal{KL}$ function. However, proving that a candidate $\mathcal{KL}$ function is an upper bound over the norms of state has to be done for all integers which can be difficult. On the other hand, Lyapunov functions are functional certificates for the system stability. Some results guarantee the existence of Lyapunov functions: a system is globally asymptotic stable if and only if there exists a Lyapunov function (see  e.g. [Kha02]). Being a Lyapunov function is a timeless property and only depends on dynamics. In few advantageous situations, Lyapunov functions can even be computed using numerical and optimization solvers (see the survey [GH15]).
Formally, the data of a peaks computation problem is decomposed into an autonomous discrete-time dynamical system and an objective function $\varphi$ over states. The dynamical system consists of a set of initial conditions $X^{\mathrm{i}n}$ and the dynamics of the system i.e. a function $T$ which maps a state to a new state. The associated peaks computation problem aims in maximizing $\varphi(T^{k}(x))$ under the constraints $x\in X^{\mathrm{i}n}$ and $k$ is a natural integer. The optimal solution of a peaks computation problem is then a couple $(x_{\rm opt},k_{\rm opt})$ where $x_{\rm opt}\in X^{\mathrm{in}}$ and $k_{\rm opt}$ is a natural integer. Note that when an integer $k$ is fixed, maximizing $\varphi(T^{k}(x))$ over $x\in X^{\mathrm{in}}$ is a standard maximization problem. The main difficulty of peaks computation problem is to compute an optimal $k_{\rm opt}$ i.e. an integer $n$ for which the supremum over $x\in X^{\mathrm{in}}$ of $\varphi(T^{n}(x))$ is the greatest. Our approach is to find an upper bound $K$ of the (greatest) integer maximizer. We impose to the upper bound $K$ to satisfy the following condition : for all $n$ greater than $K$, the supremum over $x\in X^{\mathrm{in}}$ of $\varphi(T^{n}(x))$ is smaller than the supremum of $\varphi(T^{K}(x))$. Once this integer $K$ computed, to solve the peaks computation problem can be done by solving a finite number of maximization problems. We can extract the optimal solution $(x_{\rm opt},k_{\rm opt})$ from one of those maximization problems. Thus we reduce the peaks computation problem to find the smallest possible integer $K$ satisfying the previous condition. We, then, provide a general formula to compute such integer. This formula relies on well-chosen pairs of a strictly increasing and continuous function on $[0,1]$, $h$, and a scalar $\beta\in(0,1)$. In this case, well-chosen means that the pair $(h,\beta)$ forms a sequence $(h(\beta^{k}))_{k}$ such that $(h(\beta^{k}))_{k}$ is greater than the sequence of the suprema of $\varphi(T^{k}(x))$ over $X^{\mathrm{in}}$. A well-chosen pair $(h,\beta)$ is said to be useful for our problem if $h(0)$ is strictly smaller than a supremum of $\varphi(T^{k}(x))$ over $X^{\mathrm{in}}$. We prove that there exists an optimal useful pair, in the sense of that the pair permits to find the greatest optimal integer. In practice, this pair is not computable. Therefore, we propose two other methods to construct useful pairs. First, we extend the class of $\mathcal{KL}$ functions to the class of functions which are possibly discontinuous, unsigned (on the state variable but nonnegative on the time variable), increasing on states and decreasing on time. We call this class of functions $\mathcal{KL}_{\rm gen}$. As for classical $\mathcal{KL}$ functions, we use $\mathcal{KL}_{\rm gen}$ functions as upper bounds on $\varphi(T^{k}(x))$ (for the classical case with $\mathcal{K}\mathcal{L}$, $\varphi$ is a norm). We establish an equivalence theorem where we prove that a useful pair $(h,\beta)$ can be constructed from a $\mathcal{KL}_{\rm gen}$ upper bound function and we can construct $\mathcal{KL}_{\rm gen}$ upper bounds from useful pairs $(h,\beta)$. The construction of useful pairs $(h,\beta)$ from $\mathcal{KL}_{\rm gen}$ upper bound functions relies on an extension of Sontag’s result [Son98, Prop. 7] for $\mathcal{KL}_{\rm gen}$. Finally, we propose a new concept of Lyapunov functions adapted for the peaks computation problem namely the Opt-Lyapunov functions. The word adapted means the existence of an Opt-Lyapunov function is equivalent to the existence of a useful pair $(h,\beta)$. Once again, the proofs of each implication are constructive. The construction of an Opt-Lyapunov function from a useful pair $(h,\beta)$ is inspired by Yoshizawa construction of Lyapunov functions [Yos66]. The main difference between classical Lyapunov functions and this new concept is that we allow Opt-Lyapunov functions to have several zeros and to take infinity values. Those two new possibities make irrelevant the use of classical comparison functions (see e.g. [Kel14] for a comprehensive survey). In our context, when we have access to classical Lyapunov functions, we can link continuous $\varphi$ with classical Lyapunov functions through comparison functions. Without this link, we need to develop a concept of weaker relations between $\varphi$ and an Opt-Lyapunov function. In this paper, we introduce the notion of certificate of compatibility to overcome the loss of the link.
In summary, the main contributions of the paper are the following:
  • we propose a novel approach of the peaks computation problem by bounding the number of static optimization problem to solve to obtain the optimal value;
  • we obtain a general upper bound method for bounded from above sequence based on homeomorphic deformations of geometric sequences;
  • we generalize the concept of $\mathcal{K}\mathcal{L}$ functions to $\mathcal{KL}_{\rm gen}$ adapted to peaks computation problems for which we can formulate peaks computation problems as a $\mathcal{KL}_{\rm gen}$ upper bounds existence ;
  • we introduce a new concept of Lyapunov functions well designed for peaks computation problem for which the comparison functions approach is replaced by a compatibility notion.
The paper is organized as follows. Section 2 presents the peaks computation problem addressed in this paper and the main difficulty related to the search of an integer as an optimal solution of the peaks computations problem. Section 3 recalls some basic results about the supremum of sequences and how it is related to their asymptotic behaviour. In Section 3, we also find the fundamental result about the existence of a useful pair of a strictly increasing and continuous function on $[0,1]$ and a scalar in $(0,1)$. This pair furnishes, in theory, the exact integer optimal solution. However, this pair is difficult to compute in practice. Then, in Section 4, after developing a simple result using the classical $\mathcal{K}\mathcal{L}$ class, we introduce the new class $\mathcal{KL}_{\rm gen}$. We prove that the existence of $\mathcal{KL}_{\rm gen}$ upper bounds of $\varphi(T^{k}(x))$ is equivalent to the existence of a useful pair. In Section 5, we present the concept of Opt-Lyapunov functions well designed for our maximization problem. The classical comparison functions associated to classical Lyapunov functions are replaced by a new notion called compatibility. We also prove that the existence of a compatible Opt-Lyapunov function is equivalent to the existence of a useful pair. In Section 6, we summarize the results obtained highlighting the concrete constructions between useful pairs of strictly increasing and continuous functions on $[0,1]$ and convergent geometric sequences, $\mathcal{KL}_{\rm gen}$ class and Opt-Lyapunov functions.
Notations: $\mathbb{R}$ stands for the set of reals, $\mathbb{R}_{+}$ the set of nonnegative reals, $\mathbb{R}^{*}$ the set of nonzero reals and $\mathbb{R}_{+}^{*}$ the set of strictly positive reals. The vector space of vectors of $d$ reals is denoted by $\mathbb{R}^{d}$. The set of natural integers is denoted by $\mathbb{N}$ whereas $\mathbb{N}^{*}$ denotes the set of nonzero natural integers.
2 Problem Statement And Motivated Example
2.1 Problem Statement
Let $X^{\mathrm{in}}$ be a subset of $\mathbb{R}^{d}$ and $T$ a nonzero self-map on $\mathbb{R}^{d}$. Let us consider the associated discrete-time dynamical system defined as follows:
$x_{0}\in X^{\mathrm{in}},\ x_{k+1}=T(x_{k})\enspace.$
Let us consider $\varphi:\mathbb{R}^{d}\mapsto\mathbb{R}$. We are interested in the reachable values that maximizes $\varphi$. More precisily, writing $T^{k}$ for the $k$-fold composition of $T$ or the identity when $k=0$, we are interested in solving the following maximization problem called a peaks computation problem:
Problem 1 (Peaks Computation Problem).
Maximize $\varphi(T^{k}(x))$ under the constraints $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$.
Without loss of generality, we can suppose that $\varphi(0)=0$; otherwise, we change $\varphi$ by $\varphi-\varphi(0)$ and then we add $\varphi(0)$ at the end of the optimization procedure.
To solve Problem 1 means that we have to compute the optimal value $\sup\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}},\ k\in\mathbb{N}\}$ and, if there exists, to characterize and to compute a couple of a vector $x\in X^{\mathrm{in}}$ and an integer $k$ which maximizes the value $\varphi(T^{k}(x))$. We can solve Problem 1 sequentially by solving a sufficient number of static maximization problems :
Problem Pk (Static Problems).
Maximize $\varphi(T^{k}(x))$ under the constraint $x\in X^{\mathrm{in}}$.
Some hypotheses can be made to ensure that Problem Pk has a finite optimal values and an optimal solution. For example, if $X^{\mathrm{in}}$ is compact and the function $\varphi\circ T^{k}$ is upper semicontinuous, Problem Pk has an optimal solution. About computations, some special instances of Problem Pk can be solved numerically more or less efficiently. For example, when $\varphi\circ T^{k}$ is concave and $X^{\mathrm{in}}$ is convex, the problem is said to be convex and numerous methods exist (see e.g. [BTN01, Ber15]), or when $\varphi\circ T^{k}$ is polynomial and $X^{\mathrm{in}}$ a basic semialgebraic set (see e.g. [AL11, Las15]). The main issue remains to guess the interesting $k$, the integer for which the optimal value of Problem Pk is the greatest. A direct computation of such $k$ seems to be very difficult. Hence, the method we use in this paper is to solve instances of Problem Pk for $k=0,\ldots,K$ where $K$ has to determined. The integer $K$ has the particularity of ensuring that the optimal value of Problem 1 corresponds to the optimal value of some Problem Pk where $k$ is smaller than $K$.
In this paper, we will adopt the following notations:
$\nu:=(\nu_{k})_{k\in\mathbb{N}}\text{ where }\forall\,k\in\mathbb{N},\ \nu_{k}:=\sup\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}}\}\text{ and }\nu_{\rm opt}:=\sup\{\nu_{k}:k\in\mathbb{N}\}$ (1)
Sometimes we will also use the terminlogy the supremum of $\nu$ to speak about $\nu_{\rm opt}$. The sequence $\nu$ corresponds to the sequence of optimal values of the maximization problems depicted at Problem Pk and $\nu_{\rm opt}$ is the optimal value of Problem 1. Following those notations, the main problem of the paper can be formulated mathematically as follows:
Problem 2 (Safe Number of Problem $P_{k}$ to Solve).
Find $K\in\mathbb{N}$ such that $\nu_{\rm opt}=\max\{\nu_{k}:k=0,\ldots,K\}$.
In this paper, we provide simple characterizations for which a solution $K$ of Problem 2 exists. Then we construct a formula based on pairs composed of a strictly increasing continuous functions on $[0,1]$ and a scalar in $(0,1)$. Finally, we propose constructions of those strictly increasing continuous functions on [0,1] based on extensions of classical tools on stability theory for discrete-time dynamical systems such as $\mathcal{K}\mathcal{L}$ class of functions and Lyapunov functions.
The developments made in the paper require real-valued sequences; hence the only assumption made in this paper is the following :
Assumption 1.
For all $k\in\mathbb{N},\ \nu_{k}=\sup\left\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}}\right\}$ is finite.
A simple situation where Assumption 1 holds when $X^{\mathrm{in}}$ is a bounded set, $T$ maps from bounded sets to bounded sets and $\varphi$ is upper semicontinuous.
2.2 An unstable linear system and an objective function with finite optimal value
Throughout the paper, we illustrate our method and the computations on the example presented in this subsection. The peaks computation problem can be solved more simply and we compare the results obtained with our methodology with the one obtained analytically.
Let us consider the following dynamics on $\mathbb{R}^{2}$:
$\mathbb{R}^{2}\ni(x,y)\mapsto T(x,y)=A\begin{pmatrix}x\\ y\end{pmatrix}\text{ where }A=\begin{pmatrix}1&1\\ \frac{1}{4}&1\end{pmatrix}.$
We parameterize the initial conditions set by a scalar $\mu\in(0,1/3]$ and we define:
$X^{\mathrm{in}}=([2\mu,1]\times[\mu,1])\cap\operatorname{Vect}\{(2,1)^{\intercal}\}.$
Finally, the objective function is parameterized by a positive scalar $p$:
$\mathbb{R}^{2}\ni(x,y):\mapsto\varphi(x,y):=y^{2}-x^{2}+px$
The associated peaks computation problem is the following:
$\begin{array}[]{rc}\displaystyle{\operatorname*{Max}_{(x,y),n}}&\varphi(A^{n}(x,y)^{\intercal})\\ s.t.&\left\{\begin{array}[]{l}(x,y)\in X^{\mathrm{in}}\\ n\in\mathbb{N}\end{array}\right.\end{array}$
The system is linear and the objective function is quadratic. However, we remark that $A$ admits $1/2$ and $3/2$ as eigenvalues. Therefore, $A$ is unstable and for some $x$, $\|A^{n}x\|$ goes to +$\infty$+ ∞ as $n$ goes to +$\infty$+ ∞. Therefore, we cannot use the method developed in [Adj23]. Indeed, the system does not admit a quadratic Lyapunov function. We also note that $\varphi$ is not concave nor convex.
As the matrix $A$ is diagonalizable, the powers of it are easy to compute and for all $n\in\mathbb{N}$ we have :
$A^{n}=\dfrac{1}{2^{n+1}}\begin{pmatrix}3^{n}+1&2(3^{n}-1)\\ \frac{3^{n}-1}{2}&3^{n}+1\end{pmatrix}\text{ and }\forall\,y\in\mathbb{R},\ A^{n}\begin{pmatrix}2y\\ y\end{pmatrix}=\left(\frac{3}{2}\right)^{n}\begin{pmatrix}2y\\ y\end{pmatrix}$
Moreover, the initial conditions set is actually a segment each element of it can be written as $(2y,y)^{\intercal}$ with $y\in[\mu,1/2]$. Using this, we have for all $y\in[\mu,1/2]$, for all $n\in\mathbb{N}$:
$\varphi_{n}(y):=\varphi(A^{n}(2y,y)^{\intercal})=-3\left(\dfrac{3}{2}\right)^{2n}y^{2}+2p\left(\dfrac{3}{2}\right)^{n}y$
Let $n\in\mathbb{N}$. The function $y\mapsto\varphi_{n}(y)$ is concave and infinitely derivable. It admits an unique maximizer on $\mathbb{R}$ which is $y_{n}^{*}:=\left(\frac{2}{3}\right)^{n}\frac{p}{3}$. The optimal value of $\varphi_{n}$ on $\mathbb{R}$ is $p^{2}/3$. The function $\varphi_{n}$ is strictly increasing before $y_{n}^{*}$ and strictly decreasing after. As we maximize on $[\mu,1/2]$, we have to check whether $y_{n}^{*}$ belongs to this interval. To avoid the case where $\frac{1}{2}$ is not used as a maximizer we assume that $p>3/2$. In this case, $y_{0}^{*}>1/2$.
As the sequence $(y_{n}^{*})_{n\in\mathbb{N}}$ strictly decreases to zero, we can define two integers:
$\underline{n}:=\lfloor\ln(3/2p)/\ln(2/3)\rfloor+1\text{ and }\overline{n}:=\lfloor\ln(3\mu/p)/\ln(2/3)\rfloor+1$
The integer $\underline{n}$ corresponds to the smallest integer such that $y_{n}^{*}<1/2$ whereas $\overline{n}$ is the smallest integer such that $y_{n}^{*}<\mu$. Moreover, as we assume that $\mu\in(0,1/3]$, we have $(\ln(3\mu/p)-\ln(3/2p))/\ln(2/3)=-(\ln(\mu)+\ln(2))\ln(3/2)\geq-(\ln(1/3)+\ln(2))/\ln(3/2)=1$. We conclude that $\overline{n}>\underline{n}$, as for all $x,y\in\mathbb{R}_{+}$ such that $x-y\geq 1$, we have $\lfloor x\rfloor\geq\lfloor y\rfloor+1$.
Finally using $\underline{n}$ and $\overline{n}$ , we can define the sequence of interest $(\nu_{n})_{n\in\mathbb{N}}$ and for all $n\in\mathbb{N}$:
$\nu_{n}=\left\{\begin{array}[]{lr}\varphi_{n}(1/2)&\text{ if }n\in[0,\underline{n}-1]\cap\mathbb{N}\\ \dfrac{p^{2}}{3}&\text{ if }n\in[\underline{n},\overline{n}-1]\cap\mathbb{N}\\ \varphi_{n}(\mu)&\text{ if }n\geq\overline{n}\end{array}\right.$ (2)
From now on, when we refer to this example, we will assume that $\mu\in(0,1/3]$ and $p\geq 3/2$.
3 Suprema and Rank Maximizer of General Sequences
3.1 Notations and Basic Results
We introduce the set of real sequences which are bounded from above:
$\Lambda:=\{u=(u_{0},u_{1},\ldots)\in\mathbb{R}^{\mathbb{N}}:\sup_{n\in\mathbb{N}}u_{n}<+\infty\}.$
For an element of $\Lambda$, we are interested in computing the supremum of its terms.
Proposition 1.
$u\in\Lambda\iff\displaystyle{\limsup_{n\to+\infty}u_{n}\in\mathbb{R}\cup\{-\infty\}}$.
For $u\in\Lambda$, we are also looking for the set of maximizers i.e. $\operatorname{Argmax}(u):=\{k\in\mathbb{N}:u_{k}=\displaystyle{\sup_{n\in\mathbb{N}}u_{n}}\}$. For computations purpose, we need to characterize $\operatorname{Argmax}(u)$. Consequently, we introduce, for $u\in\Lambda$, the two sets of ranks:
$\Delta_{u}:=\left\{k\in\mathbb{N}:\max_{0\leq j\leq k}u_{j}\geq\sup_{j>k}u_{j}\right\}\ \text{ and }\ \Delta^{s}_{u}:=\left\{k\in\mathbb{N}:\max_{0\leq j\leq k}u_{j}>\sup_{j>k}u_{j}\right\}$
It is easy to see that $\Delta_{u}\subseteq\Delta^{s}_{u}$. Moreover, if $k$ belongs to $\Delta_{u}$ (resp. $\Delta^{s}_{u}$), then any integer greater than $k$ belongs to $\Delta_{u}$ (resp. $\Delta^{s}_{u}$). Besides, we will need the set of ranks for which the associated term is greater than the limsup (resp. strictly greater):
${\mathrm{R}}_{u}:=\left\{k\in\mathbb{N}:u_{k}\geq\limsup_{n\to+\infty}u_{n}\right\}\ \text{ and }\ {\mathrm{R}}_{u}^{s}:=\left\{k\in\mathbb{N}:u_{k}>\limsup_{n\to+\infty}u_{n}\right\}$
It is obvious that ${\mathrm{R}}_{u}^{s}\subseteq{\mathrm{R}}_{u}$. Note that even if $u\in\Lambda$, ${\mathrm{R}}_{u}$ and hence ${\mathrm{R}}_{u}^{s}$ can be empty. To construct the characterizations, we need to consider the smallest elements of $\Delta_{u}$, $\Delta^{s}_{u}$, ${\mathrm{R}}_{u}$ and ${\mathrm{R}}_{u}^{s}$.
$\mathbf{K}_{u}=\inf\Delta_{u};\ \qquad\mathbf{K}^{s}_{u}=\inf\Delta^{s}_{u};\ \qquad k_{u}=\inf{\mathrm{R}}_{u}\quad\text{ and }\ k^{s}_{u}=\inf{\mathrm{R}}_{u}^{s}$
We recall that the infimum of the empty set is equal to +$\infty$+ ∞. Hence, $\mathbf{K}_{u}<+\infty$ (resp. $\mathbf{K}^{s}_{u}<+\infty$) if and only if $\Delta_{u}\neq\emptyset$ (resp. $\Delta^{s}_{u}\neq\emptyset$); and $k_{u}<+\infty$ (resp. $k^{s}_{u}<+\infty$) if and only if ${\mathrm{R}}_{u}\neq\emptyset$ (resp. ${\mathrm{R}}_{u}^{s}\neq\emptyset$). Furthermore, if the sets of integers defined above are nonempty, they contain their infimum i.e. $\mathbf{K}_{u}\in\Delta_{u}$, $\mathbf{K}^{s}_{u}\in\Delta^{s}_{u}$, $k_{u}\in{\mathrm{R}}_{u}$ and $k^{s}_{u}\in{\mathrm{R}}_{u}^{s}$ when those integers are finite.
Proposition 2.
If $u\notin\Lambda$ then $\Delta_{u}=\emptyset$ and ${\mathrm{R}}_{u}=\emptyset$. Obviously, it also implies that $\Delta^{s}_{u}=\emptyset$ and ${\mathrm{R}}_{u}^{s}=\emptyset$.
Proposition 3.
Let $u\in\Lambda$. The following assertions hold:
  1. 1.
    $\Delta^{s}_{u}\neq\emptyset\iff{\mathrm{R}}_{u}^{s}\neq\emptyset$;
  2. 2.
    $\Delta_{u}\neq\emptyset\iff{\mathrm{R}}_{u}\neq\emptyset$;
  3. 3.
    $\Delta^{s}_{u}=\emptyset\iff\sup_{n\in\mathbb{N}}u_{n}=\limsup_{n\to+\infty}u_{n}$.
Proposition 4.
The following inequalities hold:
$k_{u}\leq k^{s}_{u};\qquad k_{u}\leq\mathbf{K}_{u};\qquad k^{s}_{u}\leq\mathbf{K}^{s}_{u}\quad\text{ and }\mathbf{K}_{u}\leq\mathbf{K}^{s}_{u}\enspace.$
Moreover, if $k^{s}_{u}<+\infty$, then $k^{s}_{u}\leq\mathbf{K}_{u}$.
Proposition 5.
Let $u\in\Lambda$. The following assertions hold:
  1. 1.
    $\operatorname{Argmax}(u)\subseteq\Delta_{u}$;
  2. 2.
    $\operatorname{Argmax}(u)\neq\emptyset\iff\Delta_{u}\neq\emptyset$;
  3. 3.
    $\mathbf{K}_{u}=\inf\operatorname{Argmax}(u)$;
  4. 4.
    $\operatorname{Argmax}(u)=\emptyset$ or $\operatorname{Argmax}(u)$ nonempty and unbounded $\implies\sup_{n\in\mathbb{N}}u_{n}=\limsup_{k\to+\infty}u_{k}$;
  5. 5.
    If $\Delta^{s}_{u}\neq\emptyset$ then $\mathbf{K}^{s}_{u}=\max\operatorname{Argmax}(u)$.
Note that $\operatorname{Argmax}(u)$ can be nonempty and bounded when $\Delta^{s}_{u}$ is empty. In this case, by the third statement of Prop 3, we have $\sup_{n\in\mathbb{N}}u_{n}=\limsup_{n\to+\infty}u_{n}$. However, the sequence finitely touches $\limsup_{n\to+\infty}u_{n}$ before approaching, up to a subsequence, $\limsup_{n\to+\infty}u_{n}$ from below (strictly).
Example 1.
Let us consider the peaks computation problem depicted at Subsection 2.2. The sequence $(\nu_{k})_{k\in\mathbb{N}}$ presented at Eq. (2) has a greatest maximizer as $\limsup_{k\to+\infty}\nu_{k}=\lim_{k\to+\infty}\varphi_{k}(\mu)=-\infty$. The set $\operatorname{Argmax}(u)$ would not be bounded from above if $\mu=0$ and thus $\overline{n}=+\infty$. Indeed, in this case, we would get $\limsup_{k\to+\infty}\nu_{k}=p^{2}/3=\sup_{k\in\mathbb{N}}\nu_{k}$. ∎
Proposition 6.
Let $u\in\Lambda$. The following assertions hold:
  1. 1.
    Assume that ${\mathrm{R}}_{u}\neq\emptyset$. Let $k\geq k_{u}$ such that $\sup_{k_{u}\leq j\leq k}u_{j}\geq\sup_{j>k}u_{j}$ then $k\in\Delta_{u}$;
  2. 2.
    Assume that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. Let $k\geq k^{s}_{u}$ such that $\sup_{k^{s}_{u}\leq j\leq k}u_{j}\geq\sup_{j>k}u_{j}$ then $k\in\Delta_{u}$;
  3. 3.
    Assume that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. For all $\mathbf{K}_{u}\leq k$,
    $\sup_{0\leq j\leq k}u_{j}=\sup_{k^{s}_{u}\leq j\leq k}u_{j}=\sup_{k_{u}\leq j\leq k}u_{j}=\sup_{j\in\mathbb{N}}u_{j}\enspace.$
In summary, for $u\in\Lambda$, to compute $\sup_{k\in\mathbb{N}}u_{k}$, we need to study first the emptiness of ${\mathrm{R}}_{u}^{s}$. Indeed, if ${\mathrm{R}}_{u}^{s}$ is empty, we know that (the third statement of Prop. 3) $\sup_{n\in\mathbb{N}}u_{n}$ is equal to $\limsup_{n}u_{n}$. If ${\mathrm{R}}_{u}^{s}$ is not empty, then we have to compute $\mathbf{K}_{u}$ to find $\sup_{k\in\mathbb{N}}u_{k}$ and a maximizer (the second statement of Prop. 5). However, to identify $\mathbf{K}_{u}$, we need knowledge on the past and the future of the sequence. The good point is that any over-approximation $k$ of $\mathbf{K}_{u}$ permits to know $\sup_{k\in\mathbb{N}}u_{k}$ by computing $\sup_{0\leq j\leq k}u_{j}$ (the third statement of Prop 6).
3.2 A Formula Based on Pairs of Strictly Increasing Continuous Functions and Convergent Geometric Sequences
Let us write $\mathbb{I}\left(\mathbb{R}\right)$ the set of all intervals (convex sets) on $\mathbb{R}$ (open, closed, or semi-open). For a function $f$ from $\mathbb{R}$ to $\mathbb{R}$, we define the set of intervals:
$\mathrm{SIC}(f)=\{I\in\mathbb{I}\left(\mathbb{R}\right):f:I\mapsto f(I)\text{ is strictly increasing and continuous }\}.$
Recall that a strictly increasing function $f$ is continuous on an interval $[a,b]$ if and only if $f$ is surjective from $[a;b]$ to $[f(a);f(b)]$. Hence, if $I$ belongs to $\mathrm{SIC}(f)$, then $f$ is bijective from $I$ to $f(I)$ and we can define its inverse on $I$ i.e. a function $g:f(I)\mapsto I$ such that $f\circ g=g\circ f=\mathrm{Id}$ where $\mathrm{Id}:x\mapsto x$ (viewed as a function from $I$ to itself). For $I\in\mathrm{SIC}(f)$, we denote by ${{f}_{\left(I\right)}^{-1}}$ the inverse of $f$ on $I$. Note that ${{f}_{\left(I\right)}^{-1}}$ is also strictly increasing and continuous (surjective) from $f(I)$ to $I$. Finally, we pay a particular attention to functions strictly increasing and continuous on $[0,1]$ and we write :
$\Omega([0,1]):=\{h:\mathbb{R}\mapsto\mathbb{R}:[0,1]\in\mathrm{SIC}(h)\}.$
The functions of $\Omega([0,1])$ are part of the upper bound on the sequence of interest $u$. To complete the upper bound, we need a convergent geometric sequence. We introduce, for $u\in\mathbb{R}^{\mathbb{N}}$ the following sets:
$\mathbb{\Gamma}(u):=\{(h,\beta)\in\Omega([0,1])\times(0,1):u_{k}\leq h(\beta^{k}),\ \forall\,k\in\mathbb{N}\}$
and
$\mathbb{\Gamma}_{\rm fun}(u):=\{h\in\Omega([0,1]):\exists\,\beta\in(0,1)\text{s.t. }(h,\beta)\in\mathbb{\Gamma}(u)\}\enspace.$
Finally, we introduce for $u\in\mathbb{R}^{\mathbb{N}}$ and $h\in\mathbb{\Gamma}_{\rm fun}(u)$:
$\mathbb{\Gamma}_{\rm sc}(u,h):=\{\beta\in(0,1):(h,\beta)\in\mathbb{\Gamma}(u)\}\enspace.$
[…]
Proposition 7.
Let $u\in\Lambda$. Then :
  1. 1.
    For all $h\in\mathbb{\Gamma}_{\rm fun}(u)$, we have $\limsup_{k\to+\infty}u_{k}\leq h(0)$;
  2. 2.
    For all $k\in\mathbb{N}^{*}$, $u_{k}<h(1)$ and $u_{0}\leq h(1)$.
In Prop. 7, the first statement only requires $h$ to be upper semicontinuous. However, a strictly increasing function $f$ from $I$ to $f(I)$ is surjective if and only if it is continuous on $I$.
Corollary 1.
Let $u\in\Lambda$. If there exists $h\in\mathbb{\Gamma}_{\rm fun}(u)$ such that $u_{0}=h(1)$ then $u_{0}=\max_{k\in\mathbb{N}}u_{k}$.
It is easy to find a pair of a strictly increasing continuous function and a convergent geometric sequence to construct an upper bound of the sequence $u$ when the supremum of $u$ is finite.
Proposition 8.
$u\in\Lambda$ if and only if $\mathbb{\Gamma}(u)\neq\emptyset$.
Proof.
Suppose $u\in\Lambda$. By definition, $\sup_{n\in\mathbb{N}}u_{n}\in\mathbb{R}$ and the function $h:x\mapsto x+\sup_{n\in\mathbb{N}}u_{n}$ belongs to $\Omega([0,1])$ and for all $\beta\in(0,1)$ and all $k\in\mathbb{N}$, $u_{k}\leq\beta^{k}+\sup_{n\in\mathbb{N}}u_{n}$ and $\mathbb{\Gamma}(u)\neq\emptyset$. Now, let $u\in\mathbb{R}^{\mathbb{N}}$ and suppose that $\mathbb{\Gamma}(u)\neq\emptyset$. Let $(h,\beta)\in\mathbb{\Gamma}(u)$. From Prop. 7 and as $h$ is finite-valued then $\limsup_{k}u_{k}\leq h(0)\in\mathbb{R}$. From Prop. 1, this implies that $u\in\Lambda$. ∎
Some upper bounds $h(\beta^{k})$ appearing in $\mathbb{\Gamma}(u)$ cannot be used in our formula. It depends on how $h(0)$ is placed with respect to the terms of $u$.
Definition 1 (Useful Strictly Increasing Continuous Functions).
Let $u\in\Lambda$ such that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. A function $h\in\mathbb{\Gamma}_{\rm fun}(u)$ is said to be useful for $u$ if the set :
$\mathcal{S}(u,h):=\{k\in\mathbb{N}:u_{k}>h(0)\}$ (3)
is nonempty.
By extension, a pair $(h,\beta)\in\mathbb{\Gamma}(u)$ is useful for $u$ if $h$ is useful for $u$.
From the first statement of Prop. 7, if $h$ is useful for $u$, there exists $k\in\mathbb{N}$ such that $u_{k}>h(0)\geq\limsup_{n\to+\infty}u_{n}$ i.e. ${\mathrm{R}}_{u}^{s}\neq\emptyset$. We will see that the contrary also holds for i.e. if ${\mathrm{R}}_{u}^{s}\neq\emptyset$, then there exists $h\in\mathbb{\Gamma}_{\rm fun}(u)$ useful for $u$.
In Prop. 9, we present some theoretical properties on $\mathbb{\Gamma}_{\rm fun}(u)$ and $\mathbb{\Gamma}_{\rm sc}(u,h)$.
Proposition 9.
Let $u\in\Lambda$.
  1. 1.
    Let $h\in\mathbb{\Gamma}_{\rm fun}(u)$. Let us define :
    $\underline{\beta}_{u,h}:=\left\{\begin{array}[]{lr}\displaystyle{\sup_{k\in\mathcal{S}(u,h)\cap\mathbb{N}^{*}}\exp\left(\dfrac{\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))}{k}\right)}&\text{ if }\mathcal{S}(u,h)\cap\mathbb{N}^{*}\neq\emptyset\\ \\ 0&\text{ otherwise}\end{array}\right.;$
    Then $\mathbb{\Gamma}_{\rm sc}(u,h)$ is the interval $(0,1)$ if $\underline{\beta}_{u,h}=0$ and $[\underline{\beta}_{u,h},1)$ if $\underline{\beta}_{u,h}>0$.
  2. 2.
    For all finite family families $\{(h_{1},\beta_{1}),(h_{2},\beta_{2}),\ldots,(h_{n},\beta_{n})\}\subset\mathbb{\Gamma}(u)$:
    $\max_{i=1,\ldots,n}\beta_{i}\in\bigcap_{i=1,\ldots,n}\mathbb{\Gamma}_{\rm sc}(u,h_{i});$
  3. 3.
    The set $\mathbb{\Gamma}_{\rm fun}(u)$ is a convex sublattice (for the classical order, $\min$ and $\max$ on functions) subset of $\Omega([0,1])$ i.e.
    • (lattice) for all finite families $\{h_{1},h_{2},\ldots,h_{n}\}\subset\mathbb{\Gamma}_{\rm fun}(u)$, $\min_{i=1,\ldots,n}h_{i}\in\mathbb{\Gamma}_{\rm fun}(u)$ and $\max_{i=1,\ldots,n}h_{i}\in\mathbb{\Gamma}_{\rm fun}(u)$;
    • (convex) for all $t\in[0,1]$ and $h,g\in\mathbb{\Gamma}_{\rm fun}(u)$, $th+(1-t)g\in\mathbb{\Gamma}_{\rm fun}(u)$.
  4. 4.
    For all finite families $\{h_{1},h_{2},\ldots,h_{n}\}\subset\mathbb{\Gamma}_{\rm fun}(u)$, we have actually:
    $\mathbb{\Gamma}_{\rm sc}(u,\min_{i=1,\ldots,n}h_{i})=\bigcap_{i=1,\ldots,n}\mathbb{\Gamma}_{\rm sc}(u,h_{i})\subseteq\mathbb{\Gamma}_{\rm sc}(u,\max_{i=1,\ldots,n}h_{i})$
Proof.
1. Let $h\in\mathbb{\Gamma}_{\rm fun}(u)$. As $h$ is strictly increasing on $[0,1]$, it follows that if $\beta\in\mathbb{\Gamma}_{\rm sc}(u,h)$ it also holds for all $\beta^{\prime}\in[\beta,1)$. Now, it suffices to prove that $\underline{\beta}_{u,h}$ is the infimum of $\mathbb{\Gamma}_{\rm sc}(u,h)$ and it belongs to $\mathbb{\Gamma}_{\rm sc}(u,h)$ when it is strictly positive. First, we suppose that $\underline{\beta}_{u,h}=0$. By definition, this is the same as for all $k\in\mathbb{N}^{*}$, $u_{k}\leq h(0)$. As $h$ is strictly increasing, we have for all $\beta\in(0,1)$, for all $k\in\mathbb{N}^{*}$, $u_{k}\leq h(0)<h(\beta^{k})$. We have also $u_{0}\leq h(1)=h(\beta^{0})$ for all $\beta\in(0,1)$ as $h\in\mathbb{\Gamma}_{\rm fun}(u)$. Then we have $\mathbb{\Gamma}_{\rm sc}(u,h)=(0,1)$ and $\underline{\beta}_{u,h}=0=\inf\mathbb{\Gamma}_{\rm sc}(u,h)$. Now, suppose that $\underline{\beta}_{u,h}>0$. If $k=0$, $u_{0}\leq h(\underline{\beta}_{u,h}^{0})$ as $h\in\mathbb{\Gamma}_{\rm fun}(u)$. If $k\notin\mathcal{S}(u,h)$ and $k\neq 0$ then $u_{k}\leq h(0)<h(\underline{\beta}_{u,h}^{k})$. Let $k\in\mathcal{S}(u,h)\cap\mathbb{N}^{*}$, $\exp(\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))/k)\leq\underline{\beta}_{u,h}$ and thus ${{h}_{\left([0,1]\right)}^{-1}}(u_{k})\leq\underline{\beta}_{u,h}^{k}$ and then $u_{k}=h(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))\leq h(\underline{\beta}_{u,h}^{k})$. Let consider $\beta\in\mathbb{\Gamma}_{\rm sc}(u,h)$. We have for all $k\in\mathbb{N}$, $u_{k}\leq h(\beta^{k})$. Let $k\in\mathcal{S}(u,h)\cap\mathbb{N}^{*}$. By simple calculus, we have $\exp(\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))/k)\leq\beta$. Taking the supremum over $\mathcal{S}(u,h)\cap\mathbb{N}^{*}$, we obtain $\underline{\beta}_{u,h}\leq\beta<1$. We have proved that $\underline{\beta}_{u,h}\in(0,1)$, for all $k\in\mathbb{N}$, $u_{k}\leq h(\underline{\beta}_{u,h}^{k})$ and for all $\beta\in\mathbb{\Gamma}_{\rm sc}(u,h)$, $\underline{\beta}_{u,h}\leq\beta$ We conclude that $\underline{\beta}_{u,h}\in\mathbb{\Gamma}_{\rm sc}(u,h)$ and $\underline{\beta}_{u,h}=\inf\mathbb{\Gamma}_{\rm sc}(u,h)$.
2. The result follows from the fact that all elements of $\mathbb{\Gamma}_{\rm fun}(u)$ and for all $k\in\mathbb{N}^{*}$, the functions $x\mapsto x^{k}$ are increasing on the positive reals set.
3. (Lattice) First, the set $\Omega([0,1])$ is a lattice as finitely $\min$ or $\max$ of strictly increasing continuous functions are still strictly increasing and continuous. Now, it suffices to prove the results for the case $n=2$. Let $h_{1},h_{2}\in\mathbb{\Gamma}_{\rm fun}(u)$. There exist $\beta_{1},\beta_{2}\in(0,1)$ such that for all $k\in\mathbb{N}$, $u_{k}\leq h_{1}(\beta_{1}^{k})$ and $u_{k}\leq h_{2}(\beta_{2}^{k})$. Let us write $\beta=\max\{\beta_{1},\beta_{2}\}$. Then for all $k\in\mathbb{N}$, $u_{k}\leq\min\{h_{1}(\beta_{1}^{k}),h_{2}(\beta_{2}^{k})\}\leq\min\{h_{1}(\beta^{k}),h_{2}(\beta^{k}\}=\min\{h_{1},h_{2}\}(\beta^{k})$. The same techniques can be used for the maximum.
(Convexity) Let $t\in[0,1]$, $h_{1},h_{2}\in\mathbb{\Gamma}_{\rm fun}(u)$. There exist $\beta_{1},\beta_{2}\in(0,1)$ such that for all $k\in\mathbb{N}$, $u_{k}\leq h_{1}(\beta_{1}^{k})$ and $u_{k}\leq h_{2}(\beta_{2}^{k})$. Then $u_{k}=tu_{k}+(1-t)u_{k}\leq th_{1}(\beta_{1}^{k})+(1-t)h_{2}(\beta_{2}^{k})\leq th_{1}(\max\{\beta_{1},\beta_{2}\}^{k})+(1-t)h_{2}(\max\{\beta_{1},\beta_{2}\}^{k})=(th_{1}+(1-t)h_{2})(\max\{\beta_{1},\beta_{2}\}^{k})$. We conclude that $th_{1}+(1-t)h_{2}\in\mathbb{\Gamma}_{\rm fun}(u)$.
4. Let $\beta\in\mathbb{\Gamma}_{\rm sc}(u,\min_{i=1,\ldots,n}h_{i})$. Then $u_{k}\leq\min_{i=1,\ldots,n}h_{i}(\beta^{k})\leq h_{i}(\beta^{k})$ for all $k\in\mathbb{N}$. Hence, $\beta\in\cap_{i=1,\ldots,n}\mathbb{\Gamma}_{\rm sc}(u,h_{i})$. Now, let $\beta\in\cap_{i=1,\ldots,n}\mathbb{\Gamma}_{\rm sc}(u,h_{i})$. For all $i=1,\ldots,n$, for all $k\in\mathbb{N}$, $u_{k}\leq h_{i}(\beta^{k})$. Taking the minimum over $i$ leads to the conclusion. The inclusion on the right-hand side follows readily from the definition of the finite maximum of functions. ∎
We present the main result of this section. It asserts that we can find an optimal useful affine function for $u$ when ${\mathrm{R}}_{u}^{s}\neq\emptyset$.
Theorem 1 (Existence of a Useful Optimal Affine Function).
Let $u\in\Lambda$ such that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. Then, there exist $a>0$, $b\in(0,1)$ and $c\in\mathbb{R}$ such that:
  1. 1.
    $\{k\in\mathbb{N}:u_{k}>c\}\neq\emptyset$;
  2. 2.
    $u_{k}\leq ab^{k}+c$ for all $k\in\mathbb{N}$;
  3. 3.
    $u_{\mathbf{K}^{s}_{u}}=ab^{\mathbf{K}^{s}_{u}}+c$.
Proof.
If ${\mathrm{R}}_{u}^{s}$ is nonempty then $u_{\mathbf{K}^{s}_{u}}=\max_{k\in\mathbb{N}}u_{k}>\limsup_{k\to\infty}u_{k}\in\mathbb{R}\cup\{-\infty\}$ and for all $t\in(\limsup_{k\to\infty}u_{k},\max_{k\in\mathbb{N}}u_{k})$ there exists $k\in\mathbb{N}$, such that $u_{k}>t$. So let $c$ be in $(\limsup_{k\to\infty}u_{k},\sup_{k\in\mathbb{N}}u_{k})$. Now as $c>\limsup_{k\to+\infty}u_{k}$, there exists $N_{c}\in\mathbb{N}$ s.t. $\sup_{k\geq N_{c}}x_{k}\leq c$. This is the same as for all $k\geq N_{c}$, $u_{k}\leq c$. We conclude that for all $k\geq N_{c}$, $u_{k}\leq c+ab^{k}$ for all $a>0$ and all $b\in(0,1)$. Note that as $c<\max_{k\in\mathbb{N}}u_{k}$ and $\mathbf{K}^{s}_{u}$ is the greatest element of the argmax of $u$ then $N_{c}$ is strictly greater than $\mathbf{K}^{s}_{u}$. Now let introduce the set :
$S_{c}:=\{(a,b)\in\mathbb{R}_{+}^{*}\times(0,1):u_{\mathbf{K}^{s}_{u}}=\max_{k\in\mathbb{N}}u_{k}=ab^{\mathbf{K}^{s}_{u}}+c\}.$
It is easy to see that $S_{c}$ is nonempty, indeed, for all $b$ in $(0,1)$, $(\frac{u_{\mathbf{K}^{s}_{u}}-c}{b^{\mathbf{K}^{s}_{u}}},b)\in S_{c}$. Moreover, if $\mathbf{K}^{s}_{u}>0$, for all $(a,b)\in S_{c}$, for all $0<k<\mathbf{K}^{s}_{u}$, $u_{k}\leq u_{\mathbf{K}^{s}_{u}}=ab^{\mathbf{K}^{s}_{u}}+c<ab^{k}+c$ as $b\in(0,1)$. We also have $u_{0}\leq u_{\mathbf{K}^{s}_{u}}=ab^{\mathbf{K}^{s}_{u}}+c\leq a+c$ as $b\in(0,1)$. Hence, the result $u_{k}\leq ab^{k}+c$ holds on the sets $[0,\mathbf{K}^{s}_{u}]\cap\mathbb{N}$ and $[N_{c},+\infty)\cap\mathbb{N}$ with any element $(a,b)\in S_{c}$. It suffices to find $(a,b)\in S_{c}$, such that $u_{k}\leq ab^{k}+c$ for all $k\in(\mathbf{K}^{s}_{u},N_{c}]\cap\mathbb{N}$. Recall that we have $\mathbf{K}^{s}_{u}+1\leq N_{c}$.
Let us write $(-\infty,0)\ni\gamma:=\max_{k\in(\mathbf{K}^{s}_{u},N_{c}]}(u_{\mathbf{K}^{s}_{u}}-u_{k})/(\mathbf{K}^{s}_{u}-k)$. This cannot be zero since $\mathbf{K}^{s}_{u}$ is the maximal integer $k$ such that $u_{k}=\max_{n\in\mathbb{N}}u_{n}$. Now, let $(a,b)\in\mathbb{R}_{+}\times(0,1)$ and define $f_{a,b}:\mathbb{R}\ni y\mapsto ab^{y}+c$. The function $f_{a,b}$ is derivable and strictly convex and $f^{\prime}_{a,b}(y)=ab^{y}\ln(b)$ for all $y\in\mathbb{R}$. If $f^{\prime}_{a,b}(\mathbf{K}^{s}_{u})=\gamma$ and $(a,b)\in S_{c}$ then, by convexity of $f_{a,b}$, the tangent of $f$ at $y=\mathbf{K}^{s}_{u}$ satisfies for all $k\in\mathbb{N}$, $f(k)\geq f^{\prime}(\mathbf{K}^{s}_{u})(k-\mathbf{K}^{s}_{u})+f(\mathbf{K}^{s}_{u})$ which is the same as $ab^{k}+c\geq\gamma(k-\mathbf{K}^{s}_{u})+u_{\mathbf{K}^{s}_{u}}$. By definition of $\gamma$, we have for all $k\in(\mathbf{K}^{s}_{u},N_{c})$, $(\mathbf{K}^{s}_{u}-k)\gamma\leq u_{\mathbf{K}^{s}_{u}}-u_{k}$ and so $u_{k}\leq u_{\mathbf{K}^{s}_{u}}+(k-\mathbf{K}^{s}_{u})\gamma$. We conclude that if $f^{\prime}_{a,b}(\mathbf{K}^{s}_{u})=\gamma$ and $(a,b)\in S_{c}$, $u_{k}\leq ab^{k}+c$ for all $k\in(\mathbf{K}^{s}_{u},N_{c}]$. Finally, we are looking for $(a,b)\in\mathbb{R}_{+}\times(0,1)$ such that :
$f_{a,b}(\mathbf{K}^{s}_{u})=u_{\mathbf{K}^{s}_{u}}\text{ i.e. }(a,b)\in S_{c}\text{ and }f^{\prime}_{a,b}(\mathbf{K}^{s}_{u})=\gamma$
From $(a,b)\in S_{c}$, we get $ab^{\mathbf{K}^{s}_{u}}=u_{\mathbf{K}^{s}_{u}}-c$. Hence, $f^{\prime}_{a,b}(\mathbf{K}^{s}_{u})=\gamma$ if and only if $\ln(b)=\frac{\gamma}{u_{\mathbf{K}^{s}_{u}}-c}<0$. So $b=\exp(\frac{\gamma}{u_{\mathbf{K}^{s}_{u}}-c})\in(0,1)$ and $a=(u_{\mathbf{K}^{s}_{u}}-c)\exp(-\frac{\mathbf{K}^{s}_{u}\gamma}{u_{\mathbf{K}^{s}_{u}}-c})>0$. ∎
Th 1 brings deeper details compared with Prop. 8. The existence of a useful and optimal function is important for the construction and the optimality of the proposed formula. Moreover, the proof of Th 1 exhibits the parameters of the affine function to choose. Unfortunately (but naturally) all the parameters depend on $\mathbf{K}^{s}_{u}$ and $u_{\mathbf{K}^{s}_{u}}$ what is exactly what we want to compute.
Corollary 2.
Let $u\in\Lambda$ such that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. Then, there exists a pair $(h,\beta)\in\mathbb{\Gamma}(u)$ useful for $u$ such that $u_{\mathbf{K}^{s}_{u}}=h(\beta^{\mathbf{K}^{s}_{u}})$.
Proposition 10.
Let $u\in\Lambda$. Then ${\mathrm{R}}_{u}^{s}\neq\emptyset$ if and only if there exists $h\in\mathbb{\Gamma}_{\rm fun}(u)$ useful for $u$.
Proof.
Let $u\in\Lambda$ such that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. From Th 1, there exists a function $h$ of the form $x\mapsto ax+c$ which satisfies $\mathcal{S}(u,h)\neq\emptyset$. For the reverse implication, let assume ${\mathrm{R}}_{u}^{s}=\emptyset$. From Prop. 7, for all $h\in\mathbb{\Gamma}_{\rm fun}(u)$, we have $\mathcal{S}(u,h)\subseteq{\mathrm{R}}_{u}^{s}$. This proves that the set $\mathcal{S}(u,h)$ is empty for all $h\in\mathbb{\Gamma}_{\rm fun}(u)$. ∎
Example 2.
Let us consider the peaks computation problem mentioned at Subsection 2.2. We now construct an element $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for the sequence $\nu$ presented at Eq. (2).
Let $n_{0}$ be the smallest integer $n$ such that $\varphi_{n}(\mu)<0$. With simple computations, we get $n_{0}=\lfloor\ln(2p/(3\mu))/\ln(3/2)\rfloor+1$. As $\ln(2p/(3\mu))/\ln(3/2)-\ln(3\mu/p)/\ln(2/3)=(\ln(2p)-\ln(p))/\ln(3/2)$, we have $n^{0}\geq\overline{n}$. For all $n\geq\overline{n}$, $\nu_{n}=\varphi_{n}(\mu)$ and thus for all $n\geq n_{0}$, $\nu_{n}<0$.
Following the proof of Th 1, we can impose to $h$ to be linear i.e. $h:x\mapsto ax$. Indeed, the value $c$ of Th 1 can be chosen to be 0 as we have $\lim_{k\to+\infty}\nu_{k}=-\infty$ and $p^{2}/3>0$. In this case, if $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$, $h$ is useful for $\nu$. A simple upper bound over $(\nu_{n})_{n}$ is furnished from any couple $(a,\beta)$ such that:
$a>0,\ \beta\in(0,1)\text{ and }a\beta^{n_{0}}\geq p^{2}/3$
For example, we can choose $a\beta^{n_{0}}=p^{2}/3$ for which $a=2p^{2}/3$ and $\beta=2^{-1/n_{0}}$ is a solution. ∎
Th. 1 provides a theoretical result and to have in hand an optimal function $h\in\mathbb{\Gamma}_{\rm fun}(u)$ useful for $u$ is very unlikely to happen. However, if we have any function $h\in\mathbb{\Gamma}_{\rm fun}(u)$, we are able to produce a safe upper bound of $\mathbf{K}^{s}_{u}$ and then a solution of Problem 2. To compute it, we define, for $u\in\mathbb{R}^{\mathbb{N}}$, we introduce the function $\mathfrak{F}_{u}$, from ×$\mathbb{N}\times\Omega([0,1])\times(0,1)\mapsto\mathbb{R}\cup\{+\infty\}$ defined for all $k\in\mathbb{N}$, and $h\in\Omega([0,1])$ and $\beta\in(0,1)$ by:
$\mathfrak{F}_{u}(k,h,\beta):=\left\{\begin{array}[]{lr}\dfrac{\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))}{\ln(\beta)}&\text{ if }(h,\beta)\in\mathbb{\Gamma}(u)\text{ and }k\in\mathcal{S}(u,h)\\ +\infty&\text{ otherwise}\end{array}\right.$ (4)
We will show that $\mathfrak{F}_{u}$ has the desired properties to solve Problem 2.
Proposition 11.
Let us take a pair $(h,\beta)\in\mathbb{\Gamma}(u)$ useful for $u$ and let $k\in\mathcal{S}(u,h)$. The following statements hold:
  1. 1.
    $\mathfrak{F}_{u}(k,h,\beta)$ is well-defined and strictly positive if $u_{0}<h(1)$ and null if $k=0$ and $u_{0}=h(1)$;
  2. 2.
    $\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1=\min\{j\in\mathbb{N}:h(\beta^{j})<u_{k}\}=\min\{j\in\mathbb{N}:\beta^{j}<\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k})\}$;
  3. 3.
    $\mathfrak{F}_{u}(k,h,\beta)\geq k$;
  4. 4.
    For all $j\in\mathbb{N}$, $j>\mathfrak{F}_{u}(k,h,\beta)$, $u_{j}<u_{k}$;
  5. 5.
    Let $j\in\mathcal{S}(u,h)$. If $u_{j}\leq u_{k}$ then $\mathfrak{F}_{u}(k,h,\beta)\leq\mathfrak{F}_{u}(j,h,\beta)$.
Proof.
1. Let $(h,\beta)\in\mathbb{\Gamma}(u)$. As $h\in\mathbb{\Gamma}_{\rm fun}(u)$ and we only consider $u_{k}>h(0)$, we have ${{h}_{\left([0,1]\right)}^{-1}}(u_{k})>0$. We also have by definition of $h$, if $k\in\mathbb{N}^{*}$, $u_{k}<h(1)$ meaning that ${{h}_{\left([0,1]\right)}^{-1}}(u_{k})\in(0,1)$. If $u_{0}<h(1)$, the strict inequality is also valid at $k=0$. By definition, $\beta\in(0,1)$. We can apply the natural logarithm to ${{h}_{\left([0,1]\right)}^{-1}}(u_{k})$ and $\beta$; those values are both strictly negative and then the quotient is positive. If we have $u_{0}=h(1)$ and $k=0$, $\mathfrak{F}_{u}(k,h,\beta)=0$.
2. As $h$ is strictly increasing and continuous on $[0,1]$, $\beta\in(0,1)$ and $k\in\mathcal{S}(u,h)$, there exists $j\in\mathbb{N}$ such that $h(\beta^{j})<u_{k}$. Then the sets $\{j\in\mathbb{N}:h(\beta^{j})<u_{k}\}$ and $\{j\in\mathbb{N}:\beta^{j}<\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k})\}$ are equal and nonempty. Let $j\in\mathbb{N}$ s.t. $\beta^{j}<\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k})$. Applying the natural logarithm and dividing by $\ln(\beta)<0$ leads to $j>\mathfrak{F}_{u}(k,h,\beta)$ and as $j\in\mathbb{N}$ it implies that $j\geq\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1$. By definition of integer parts and as $\beta\in(0,1)$, $\ln(\beta)(\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1)<\mathfrak{F}_{u}(k,h,\beta)\ln(\beta)=\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))$. Applying the exponential leads to $\beta^{\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1}<\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k})$ and $\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1$ belongs to $\{j\in\mathbb{N}:\beta^{j}<\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k})\}$ and is its minimum integer.
3. The pair $(h,\beta)$ belongs to $\mathbb{\Gamma}(u)$ and thus for all $k\in\mathbb{N}$, $u_{k}\leq h(\beta^{k})$ which is the same as ${{h}_{\left([0,1]\right)}^{-1}}(u_{k})\leq\beta^{k}$. Applying the natural logarithm and dividing by $\ln(\beta)$ leads to the result.
4. Let $j>\mathfrak{F}_{u}(k,h,\beta)$ then $j\geq\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1$ and, as $h$ is increasing, $h(\beta^{j})\leq h(\beta^{\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1})<u_{k}$ from the second statement. Moreover, by definition of $(h,\beta)$, we have $u_{j}\leq h(\beta^{j})$. It follows that $u_{j}<u_{k}$.
[…]
We relate some structural properties of the sets $\mathbb{\Gamma}_{\rm fun}(u)$ and $\mathbb{\Gamma}_{\rm sc}(u,h)$ depicted at Prop. 9 with the computations of $\mathfrak{F}_{u}(k,h,\beta)$.
Proposition 12.
Let $u\in\Lambda$ such that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. The following properties hold:
  1. 1.
    Let $h\in\mathbb{\Gamma}_{\rm fun}(u)$ and $k\in\mathcal{S}(u,h)$. Let $\beta,\beta^{\prime}\in\mathbb{\Gamma}_{\rm sc}(u,h)$ such that $\beta\leq\beta^{\prime}$. Then $\mathfrak{F}_{u}(k,h,\beta)\leq\mathfrak{F}_{u}(k,h,\beta^{\prime})$.
  2. 2.
    Let $g,h\in\mathbb{\Gamma}_{\rm fun}(u)$. Let $k\in\mathcal{S}(u,\min\{g,h\})$. Let $\beta\in\mathbb{\Gamma}_{\rm sc}(u,\min\{g,h\})$. Then $\mathfrak{F}_{u}(k,\min\{g,h\},\beta)\leq\min\{\mathfrak{F}_{u}(k,g,\beta),\mathfrak{F}_{u}(k,h,\beta)\}$.
Proof.
1. The function $\beta\mapsto\mathfrak{F}_{u}(k,h,\beta)$ is increasing as it is the composition of (×$(\times\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k})))\circ{\cdot}^{-1}\circ\ln$ where $\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))$ is negative.
2. First from Prop 9, $f:=\min\{g,h\}\in\mathbb{\Gamma}_{\rm fun}(u)$. Second, for all $x\in[0,1]$, $f(x)\leq g(x)$ and we can apply ${{f}_{\left([0,1]\right)}^{-1}}$ which is strictly increasing on $[0,1]$ and $x\leq\displaystyle{{f}_{\left([0,1]\right)}^{-1}}(g(x))$. Thus, we write $y=g(x)$ and thus $x=\displaystyle{{g}_{\left([0,1]\right)}^{-1}}(y)$. We conclude that ${{g}_{\left([0,1]\right)}^{-1}}(y)\leq\displaystyle{{f}_{\left([0,1]\right)}^{-1}}(y)$. Using the same technique, we can prove that ${{h}_{\left([0,1]\right)}^{-1}}(y)\leq\displaystyle{{f}_{\left([0,1]\right)}^{-1}}(y)$. The result follows from the fact that $\ln$ is increasing and $\ln(\beta)$ is negative. ∎
The first statement of Prop. 12 has to be read as follows : we have to choose the smallest possible $\beta\in\mathbb{\Gamma}_{\rm sc}(u,h)$ to reduce the approximation gap. The second statement has a similar meaning in a functional sense : if we have several functions in $\mathbb{\Gamma}_{\rm fun}(u)$ we have to consider the minimum functions to reduce the overapproximation gap.
Proposition 13.
Let $u\in\Lambda$ and $(h,\beta)\in\mathbb{\Gamma}(u)$. Then for all $j\in\mathbb{N}$, $\mathbf{K}^{s}_{u}\leq\lfloor\mathfrak{F}_{u}(j,h,\beta)\rfloor$. Moreover, if $h$ is useful for $u$, we have:
$\min_{k\in\mathcal{S}(u,h)}\mathfrak{F}_{u}(k,h,\beta)=\mathfrak{F}_{u}(\mathbf{K}^{s}_{\nu},h,\beta)\enspace.$
Proof.
By definition of $\mathfrak{F}_{u}$, if $u_{j}\leq h(0)$, the inequality holds. Now let $j\in\mathcal{S}(u,h)$, we have $u_{j}\leq\sup_{k\in\mathbb{N}}u_{k}=u_{\mathbf{K}^{s}_{u}}$. Hence, from the fifth statement of Prop. 11, $\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)\leq\mathfrak{F}_{u}(j,h,\beta)$. We conclude from the third statement of Prop. 11 that $\mathbf{K}^{s}_{u}\leq\lfloor\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)\rfloor$ and $\min_{k\in\mathcal{S}(u,h)}\mathfrak{F}_{u}(k,h,\beta)=\mathfrak{F}_{u}(\mathbf{K}^{s}_{\nu},h,\beta)$ whenever the set $\mathcal{S}(u,h)$ is nonempty. Indeed, if $\mathcal{S}(u,h)$ is nonempty, then $u_{\mathbf{K}^{s}_{u}}=\max_{k\in\mathbb{N}}u_{k}$ is also strictly greater than $h(0)$. ∎
The result of Prop. 13 is not surprising but can be directly used in a very particular context. Suppose that we have an explicit closed formula for $u_{k}$ but not derivable. The function $h$ can regularize and strictly convexify the closed formula. We thus obtain with some calculus the exact value of $\mathbf{K}^{s}_{\nu}$ as the integer part (or integer part +1) of the zero of the derivative of $s\to\mathfrak{F}_{u}(s,h,\beta)$ (extended in the reals).
We can go further than the results of Prop 13 to get a solution of Problem 2. Th. 2 constitutes the next step of Prop 13.
Theorem 2 (Solution of Problem 2).
Let $u\in\Lambda$ with ${\mathrm{R}}_{u}^{s}\neq\emptyset$. Let $h\in\mathbb{\Gamma}_{\rm fun}(u)$ be useful for $u$ and let $\beta\in\mathbb{\Gamma}_{\rm sc}(u,h)$ and $k\in\mathcal{S}(u,h)$. Then $\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor$ is a solution of Problem 2 that is :
$\max_{n\in\mathbb{N}}u_{n}=\max\{u_{k}:k=0,\ldots,\lfloor\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)\rfloor\}.$
Proof.
It suffices to combine Prop. 6 and Prop. 13. ∎
Th. 2 shows how the function depicted at Eq. (4) is used to solve Problem 2. It suffices to construct an upper bound of the sequence of interest of the form of $(h(\beta^{k}))_{k\in\mathbb{N}}$ where $h$ is strictly increasing and continuous on $[0,1]$ and $\beta\in(0,1)$; moreover, $h(0)$ has to be strictly smaller than some terms of the sequence of interest. In theory, when the limsup of the sequence of interest is not the supremum of the sequence, such couples exist and one of them is optimal (Th. 1) and $h$ can even be chosen as an affine function.
Example 3.
To give a practical view of the function provided at Eq. (4), we evaluate $\mathfrak{F}_{\nu}(k,h,\beta)$ for the sequence $\nu$ described at Eq (2) of Subsection 2.2. The pairs $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for $\nu$ considered are the ones proposed at Example 2.
We evaluate the function for different values of $p$ and $\mu$ to two choices of pairs $(h,\beta)$. In Example 2, we proposed linear functions $h$ characterized by a scalar $a$. We recall that the pair are assumed to satisfy the conditions: $a>0$, $\beta\in(0,1)$ and $a\beta^{n_{0}}\geq p^{2}/3$ where $n_{0}$ is the smallest integer $n$ such that $\varphi_{n}(\mu)<0$. First, as the function $\mathfrak{F}_{u}$ depends on the values of the term $\nu_{k}$, we show the seven first $\nu_{0},\ldots,\nu_{6}$ (rounded to the nearest hundredth) for certain values of $p$ and $\mu$ at Table 1.
$\nu_{k}$
k 0 1 2 3 4 5 6 $\mathbf{K}^{s}_{\nu}$ : $\nu_{\rm opt}$
$p=30$ $\mu=\frac{1}{3}$ 29.25 43.31 63.70 92.71 132.65 184.56 244.41 8 : 300
$\mu=\frac{1}{10}$ 29.25 43.31 63.70 92.71 132.65 184.56 244.41 11 : 300
$\mu=\frac{1}{1000}$ 29.25 43.31 63.70 92.71 132.65 184.56 244.41 22 : 300
$p=9$ $\mu=\frac{1}{3}$ 8.25 11.81 16.45 21.83 26.34 27. 25.09 5 : 27
$\mu=\frac{1}{10}$ 8.25 11.81 16.45 21.83 26.34 27. 27. 8 : 27
$\mu=\frac{1}{1000}$ 8.25 11.81 16.45 21.83 26.34 27. 27. 19 : 27
$p=3$ $\mu=\frac{1}{3}$ 2.25 2.81 3 2.95 1.58 -4.03 -20.47 2 : 3
$\mu=\frac{1}{10}$ 2.25 2.81 3 3. 3. 3. 2.94 5 : 3
$\mu=\frac{1}{1000}$ 2.25 2.81 3. 3. 3. 3. 3. 17 : 3
Table 1: Few values of the terms $\nu_{k}$ defined at Eq. (2) depending on the problem parameters $p$ and $\mu$.
In Table 1, we observe that the terms $\nu_{k}$ strictly increases until the optimal values (the smallest integer maximizer $\mathbf{K}_{\nu}$) and strictly decreases after (after the greatest integer maximizer $\mathbf{K}^{s}_{\nu}$). This observation will be formally proved and exploited later on. Another point related to this observation is that the supremum of $\nu_{k}$ for $k\notin\operatorname{Argmax}(\nu)$ seems to be either $\nu_{\underline{n}-1}$ or $\nu_{\overline{n}}$. For example, for $p=3$ and $\mu=1/3$, $\nu_{3}>\nu_{1}$ whereas for $p=9$ and $\mu=1/3$, $\nu_{4}>\nu_{5}$. Actually, we will see that the rank preceding $\underline{n}$ is not the one to consider to compute $\nu_{k}$ for $k\notin\operatorname{Argmax}(\nu)$.
Now, we evaluate the function $\mathfrak{F}_{\nu}$ for our example. To simplify, we write $\mathfrak{F}_{\nu}(k,a,\beta)$ instead of $\mathfrak{F}_{\nu}(k,h,\beta)$. We choose two pairs: first, $a=p^{2}$ and $\beta=(2/3)^{1/n_{0}}$ and second $a=2p^{2}/3$ and $\beta=(1/2)^{n_{0}}$. The values of $\lfloor\mathfrak{F}_{\nu}(k,a,\beta)\rfloor$ for these two choices are provided at Table 2.
$\lfloor\mathfrak{F}_{\nu}(k,p^{2},(2/3)^{1/n_{0}})\rfloor$ $\lfloor\mathfrak{F}_{\nu}(k,p^{2}/3,(1/2)^{1/n_{0}})\rfloor$
k $\underline{n}$ $\overline{n}$ $n_{0}$ 0 1 2 3 4 $\mathbf{K}^{s}_{\nu}$ 0 1 2 3 4 $\mathbf{K}^{s}_{\nu}$
$p$ $\mu=\frac{1}{3}$ 8 9 10 84 74 65 56 47 27 43 37 32 26 21 9
= $\mu=\frac{1}{10}$ 8 12 13 97 84 72 61 50 35 56 49 42 35 28 12
30 $\mu=\frac{1}{1000}$ 8 23 24 202 179 156 131 113 65 104 91 77 64 52 23
$p$ $\mu=\frac{1}{3}$ 5 6 7 39 33 27 22 19 18 18 15 12 9 7 7
= $\mu=\frac{1}{10}$ 5 9 10 56 47 39 32 27 27 27 21 17 13 10 9
9 $\mu=\frac{1}{1000}$ 5 20 21 118 99 82 67 58 56 56 46 36 27 21 20
$p$ $\mu=\frac{1}{3}$ 2 3 4 13 11 10 $10^{*}$ $17^{*}$ 10 5 4 3 $4^{*}$ $7^{*}$ 3
= $\mu=\frac{1}{10}$ 2 4 5 23 20 18 18 18 18 9 7 7 7 7 7
3 $\mu=\frac{1}{1000}$ 2 18 18 61 51 48 48 48 48 25 19 18 18 18 18
Table 2: Evaluations of the function $\lfloor\mathfrak{F}_{\nu}(k,h,\beta)\rfloor$ described at Eq. (4) for the sequence $\nu$ of Eq. (2) and $(h,\beta)\in\mathbb{\Gamma}(\nu)$ proposed at Ex. 2
The numbers of Table 2 at columns $\lfloor\mathfrak{F}_{\nu}(k,p^{2},(2/3)^{1/n_{0}})\rfloor$ and $\lfloor\mathfrak{F}_{\nu}(k,2p^{2}/3,(1/2)^{1/n_{0}})\rfloor$ represent safe upper bounds on the number of terms of the sequence $\nu$ of Eq. (2) to compare to obtain the supremum of the terms $\nu_{k}$. From the analysis made at Subsection 2.2, $\mathbf{K}^{s}_{\nu}$ is well identified and is equal to $\overline{n}-1$. We observe that all numbers of the concerned columns are greater than $\overline{n}-1$ as stated at Prop. 13. For both choices of $a$ and $\beta$, the best upper bound is furnished by the column $\lfloor\mathfrak{F}_{\nu}(\mathbf{K}^{s}_{\nu},a,\beta)$ as asserted at Prop. 13. Note that four integers are followed by an asterisk. In practice, those numbers are useless since not computed. Indeed, applying the fifth statement of Prop. 11, it suffices to compute a new stopping integer at $k$ when we have $\nu_{j}<\nu_{k}$ for all $j<k$. Otherwise, the stopping integer is greater than at least one previous computed integer. Moreover, for the case $a=2p^{2}/3$ and $\beta=(1/2)^{1/n_{0}}$, following Th 2, the terms $\nu_{k}$, for $k>3$, are not computed as $\mathfrak{F}_{\nu}(3,2p^{2}/3,(1/2)^{1/n_{0}})=3$ and we only have to compare $\nu_{0},\nu_{1},\nu_{2}$ and $\nu_{3}$ to obtain the maximal values of the terms $\nu_{k}$.
Finally, Table 2 highlights the importance of the choice of the pair $(h,\beta)\in\mathbb{\Gamma}(u)$ to compute a less conservative upper bound of $\mathbf{K}^{s}_{u}$. For general peaks computation problems of the form Problem 1, those upper bounds serve to fix the number of instances of Problem Pk to solve to compute the optimal value of the peaks computation problem. As Problem Pk are constrained optimization problems, a too conservative upper bound can dragstically increase the computational time. ∎
Theorem 3 (Formula Optimality).
For all $u\in\mathbb{R}^{\mathbb{N}}$, we have the following formula:
$\mathbf{K}^{s}_{u}=\inf_{(h,\beta)\in\Gamma(u)}\inf_{k\in\mathcal{S}(u,h)}\mathfrak{F}_{u}(k,h,\beta)=\inf_{(h,\beta)\in\Gamma(u)}\inf_{k\in\mathcal{S}(u,h)}\dfrac{\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(u_{k}))}{\ln(\beta)}$
and the $\inf$ can be replaced by $\min$ when ${\mathrm{R}}_{u}^{s}\neq\emptyset$.
Proof.
Let us define
$F:=\inf_{(h,\beta)\in\Gamma(u)}\inf_{k\in\mathcal{S}(u,h)}\mathfrak{F}_{u}(k,h,\beta).$
If $u\notin\Lambda$, then from Prop. 8, $\mathbb{\Gamma}(u)=\emptyset$ and thus $F=+\infty$. Moreover, from Prop. 2, $\mathbf{K}^{s}_{u}=+\infty$.
Suppose that $u\in\Lambda$. From Prop. 8, $\mathbb{\Gamma}(u)\neq\emptyset$. Assume that $\Delta^{s}_{u}$ is empty. This is the same as $\mathbf{K}^{s}_{u}=+\infty$. From Prop. 3, ${\mathrm{R}}_{u}^{s}=\emptyset$ and from Prop. 10, $\mathcal{S}(u,h)=\emptyset$ which leads to $F=+\infty$. Assume that $\Delta^{s}_{u}$ is nonempty and thus ${\mathrm{R}}_{u}^{s}\neq\emptyset$. From Prop. 13, $\mathbf{K}^{s}_{u}\leq F$. Moreover, from Corollary 2, there exists $(h,\beta)\in\mathbb{\Gamma}(u)$ such that $u_{\mathbf{K}^{s}_{u}}=h(\beta^{\mathbf{K}^{s}_{u}})>h(0)$ and $F<+\infty$. Taking the inverse of $h$ on $[0,1]$ and the natural logarithm, we get $\mathbf{K}^{s}_{u}=\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)$. By definition of $F$, $F\leq\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)=\mathbf{K}^{s}_{u}$. We conclude that $\mathbf{K}^{s}_{u}=F$. ∎
In general, we cannot rely on pre-analysis of the system as it is done in Subsection 2.2 to solve a peaks computation problem. Thanks to Th. 2, we reduce a peaks computation problem to the construction of a pair $(h,\beta)$ useful for $\nu$. In consequence, we are faced to the following problem on the sequence $\nu$ defined at Eq (1).
Problem 3.
Construct a couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for $\nu$ when ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$.
Next sections of the paper are focused on classical tools used in stability theory of dynamical systems such as $\mathcal{K}\mathcal{L}$-stability and Lyapunov functions. We show how these tools can be adapted to solve Problem 3.
4 Extended $\mathcal{K}\mathcal{L}$ Theory for Discrete-Time Systems Maximization
In this section, we come back to the initial problem that is a peaks computation problem i.e. the maximization of a function over the reachable values of a discrete-time dynamical system. We recall that the data are the following:
  • an initial conditions set $X^{\mathrm{in}}\subseteq\mathbb{R}^{d}$;
  • a nonzero map $T:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$;
  • a function $\varphi:\mathbb{R}^{d}\mapsto\mathbb{R}$ such that $\varphi(0)=0$.
We recall that we want compute $\nu_{\rm opt}=\sup\{\nu_{k}:k\in\mathbb{N}\}$ where $\nu_{k}=\sup\{\varphi(T^{k}(x):x\in X^{\mathrm{in}}\}$. To use the results of Section 3 with the sequence $\nu=(\nu_{k})_{k\in\mathbb{N}}$, we suppose that Assumption 1 holds.
We propose to simplify the notation of the maximal value of an objective function $f:\mathbb{R}^{d}\mapsto\mathbb{R}$ on $X^{\mathrm{in}}$, we introduce the following new notation:
$\overline{f_{X^{\mathrm{in}}}}:=\sup\{f(x):x\in X^{\mathrm{in}}\}$
and the set of functions which are bounded from above on $X^{\mathrm{in}}$:
$\operatorname{Fin}\left(X^{\mathrm{in}}\right):=\{f:\mathbb{R}^{d}\mapsto\mathbb{R}\cup\{+\infty\}:\overline{f_{X^{\mathrm{in}}}}<+\infty\}$
With these notations, $\nu_{k}=\overline{(\varphi\circ T^{k})_{X^{\mathrm{in}}}}$ and by following Assumption 1, we have for all $k\in\mathbb{N}$, $\varphi\circ T^{k}\in\operatorname{Fin}\left(X^{\mathrm{in}}\right)$. Following Subsection 3.2, we have to study, first, the limsup of the sequence $\nu$ and then construct, when the limsup is finite, an element $(h,\beta)$ of $\mathbb{\Gamma}(\nu)$ i.e. $h$ is strictly increasing and continuous on $[0,1]$, $\beta\in(0,1)$ and for all $k\in\mathbb{N}$, $h(\beta^{k})$ is greater than $\nu_{k}$. We also have to guarantee that if ${\mathrm{R}}_{\nu}^{s}$ is nonempty, $h$ has to be useful for $\nu$ i.e. $\mathcal{S}(\nu,h)$ is nonempty. Without this latter condition, the function $h$ is so useless as $\mathfrak{F}_{u}(k,h,\beta)=+\infty$. In this section, we will revisit $\mathcal{KL}$ class of functions to study the limsup of $\nu$ and to construct our necessary pair $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for $\nu$.
4.1 Using the classical $\mathcal{KL}$ class of functions
To study global asymptotic stability of discrete-time systems can be done by means of comparison functions (see e.g. [GGLW14]). For a comprehensive literature about comparison functions, the interested reader can refer to  [Kel14]. Comparison functions can be used for our problem to study the limsup of the sequence $\nu$.
We start by recalling some necessary definitions about comparison functions.
Definition 2 ($\mathcal{K}$/$\mathcal{K}_{\infty}$ functions classes).
A function $\alpha:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ belongs to $\mathcal{K}$ if $\alpha$ is strictly increasing continuous and $\alpha(0)=0$. If moreover, $\lim_{x\to+\infty}\alpha(x)=+\infty$, $\alpha$ is said to belong to $\mathcal{K}_{\infty}$.
Definition 3 ($\mathcal{L}$ functions class).
A function $\sigma:\mathbb{R}_{+}\to\mathbb{R}_{+}$ belongs to $\mathcal{L}$ if it is continuous, strictly decreasing and $\lim_{x\to+\infty}\sigma(x)=0$.
Definition 4 ($\mathcal{K}\mathcal{L}$ functions class).
A function $\gamma:\mathbb{R}_{+}\times\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ belongs to $\mathcal{K}\mathcal{L}$ if for all $t\in\mathbb{R}_{+}$, $s\mapsto\gamma(s,t)$ belongs to $\mathcal{K}$ and if for all $s\in\mathbb{R}_{+}$, $t\mapsto\gamma(s,t)$ belongs to $\mathcal{L}$.
The class $\mathcal{K}$ is included in $\Omega([0,1])$ but $\mathcal{K}$ functions are strictly positive except at 0 where they are null. Hence, $h\in\mathcal{K}\cap\mathbb{\Gamma}_{\rm fun}(\nu)$ is useful for $\nu$ if at least one term $\nu_{k}$ is strictly positive. Thus, in this subsection, Assumption 2 must hold.
Assumption 2.
There exists $k\in\mathbb{N}$ and $x\in X^{\mathrm{in}}$ such that $\varphi(T^{k}(x))>0$.
We recall the result obtained by Sontag.
Lemma 1 (Proposition 7 of [Son98]).
For all $\gamma\in\mathcal{K}\mathcal{L}$, there exists $\theta_{1},\theta_{2}\in\mathcal{K}_{\infty}$ such that for all $s,t\in\mathbb{R}_{+}$, $\gamma(s,t)\leq\theta_{1}(\theta_{2}(s)e^{-t})$.
Theorem 4.
Assume that there exist a function $\gamma\in\mathcal{K}\mathcal{L}$ and a function $\psi:\mathbb{R}^{d}\mapsto\mathbb{R}_{+}$ such that $\overline{\psi_{X^{\mathrm{in}}}}>0$. We also assume that for all $x\in X^{\mathrm{in}}$ and for all $k\in\mathbb{N}$:
$\varphi(T^{k}(x))\leq\gamma(\psi(x),k)\enspace.$
Then:
  1. 1.
    $\limsup_{n\to+\infty}\nu_{n}\leq 0$ and if, moreover, Assumption 2 holds then ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$;
  2. 2.
    there exists $h\in\mathcal{K}_{\infty}$ s.t. $(h,e^{-1})\in\mathbb{\Gamma}(\nu)$ and, if, moreover, Assumption 2 holds then $h$ is useful for $\nu$.
Proof.
[…]
2. This a direct consequence of the Sontag’s result (Lemma 1 here). Indeed, there exist $\theta_{1},\theta_{2}\in\mathcal{K}_{\infty}$ such that for all $x\in X^{\mathrm{in}}$ and for all $k\in\mathbb{N}$ $\varphi(T^{k}(x))\leq\gamma(\psi(x),k)\leq\theta_{1}(\theta_{2}(\psi(x))e^{-k})$. Then $\nu_{k}\leq h(e^{-1}):=\sup_{x\in X^{\mathrm{in}}}\theta_{1}(\theta_{2}(\psi(x)e^{-k}))$. It is easy to see, as $\theta_{1}$ and $\theta_{2}$ are increasing and continuous, that $h$ can be also written $y\mapsto\theta_{1}(\theta_{2}(\overline{\psi_{X^{\mathrm{in}}}}y))$. The function $h$ is thus the compositions of $\theta_{1}$, $\theta_{2}$ and $y\mapsto\overline{\psi_{X^{\mathrm{in}}}}y$ which belong to $\mathcal{K}_{\infty}$ (the third as $\overline{\psi_{X^{\mathrm{in}}}}>0$) and thus belongs to $\mathcal{K}_{\infty}$. Finally, under Assumption 2, there exists $k\in\mathbb{N}$, such that $\nu_{k}>0=h(0)$. ∎
Usually i.e. when $\varphi=\|\cdot\|$, the inequality $\varphi(T^{k}(x))\leq\gamma(\psi(x),k)$ (for $\psi=\|\cdot\|$) is true for all $x\in\mathbb{R}^{d}$ and all $k\in\mathbb{N}$ if and only if the system $x_{k+1}=T(x_{k}),x_{0}\in\mathbb{R}^{d}$ is said to be uniformly global asymptotic stability (UGAS) (see for example [NT04, Prop. 1] or [ATP10, Th. 2]. This implies that $x_{k}$ tends to 0 as $k$ tends to +$\infty$+ ∞. Thus, with a general function $\varphi$, we obtain a one-sided information on the asymptotic behaviour of the dynamics. We could obtain the other side applying the theorem for $-\varphi$.
4.2 Revisiting $\mathcal{KL}$ functions : the class $\mathcal{KL}_{\rm gen}$
In this paper, we are not interested in stability of the dynamical system, we can relax the strict increase (or decrease), the continuity and the positivity (on the state variable) conditions required for $\mathcal{KL}$ functions.
Definition 5 (The Class $\mathcal{KL}_{\rm gen}$).
A function $\gamma:\mathbb{R}\times\mathbb{R}_{+}\mapsto\mathbb{R}$ belongs to $\mathcal{KL}_{\rm gen}$ if for all $s\in\mathbb{R}$, $t\mapsto\gamma(s,t)$ is increasing and for all $t\in\mathbb{R}_{+}$, $s\mapsto\gamma(s,t)$ is decreasing.
Definition 6 ($\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bounds).
A pair $(\gamma,\theta)\in\mathcal{KL}_{\rm gen}\times\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ is a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ if and only if :
$\forall\,k\in\mathbb{N},\ \forall\,x\in X^{\mathrm{in}},\ \varphi(T^{k}(x))\leq\gamma(\theta(x),k)$
Moreover, $(\gamma,\theta)$ is said to useful for $\nu$ such that ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$ if and only if the set:
$\left\{k\in{\mathrm{R}}_{\nu}^{s}:\inf_{t\geq 0}\gamma(\overline{\theta_{X^{\mathrm{in}}}},t)<\nu_{k}\right\}$
is nonempty.
Following the notations introduced earlier, we have the equivalence:
$\nu\in\Lambda\iff\sup_{k\in\mathbb{N}}\varphi\circ T^{k}\in\operatorname{Fin}\left(X^{\mathrm{in}}\right)$
From Prop. 1 we have the equivalence, we also have:
$\nu\in\Lambda\iff\limsup_{k\to+\infty}\nu_{k}=\limsup_{k\to+\infty}\sup_{x\in X^{\mathrm{in}}}\varphi(T^{k}(x))=\limsup_{k\to+\infty}\ \overline{(\varphi\circ T^{k})_{X^{\mathrm{in}}}}\in\mathbb{R}\cup\{-\infty\}$
We can establish a first result which states that the existence of $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bounds is equivalent to the fact that the sequence $\nu$ is bounded from above. This is an adaptation of classical results dealing with stability and $\mathcal{KL}$ upper bounds.
Theorem 5 ($\mathbb{\Gamma}(\nu)$ to $\mathcal{KL}_{\rm gen}$).
The sequence $\nu=(\nu_{k})_{k\in\mathbb{N}}$ belongs to $\Lambda$ if and only if there exists a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound $(\gamma,\theta)$ for $(X^{\mathrm{in}},T,\varphi)$. If, moreover, ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$ then $(\gamma,\theta)$ can be chosen to be useful for $\nu$.
Proof.
Suppose that there exists a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound $(\gamma,\theta)$ for $(X^{\mathrm{in}},T,\varphi)$. As $s\mapsto\gamma(s,t)$ is increasing for all $t\geq 0$, we have for all $x\in X^{\mathrm{in}}$, for all $k\in\mathbb{N}$, $\gamma(\theta(x),k)\leq\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)$ and thus $\nu_{k}=\sup\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}}\}\leq\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)$ for all $k\in\mathbb{N}$. Now, taking the limsup on both sides leads to $\limsup_{k\mapsto+\infty}\nu_{k}\leq\inf_{k\in\mathbb{N}}\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)\leq\gamma(\overline{\theta_{X^{\mathrm{in}}}},0)<+\infty$ as $t\mapsto\gamma(s,t)$ is decreasing for all $s\in\mathbb{R}$.
Suppose that $\nu=(\nu_{k})_{k\in\mathbb{N}}\in\Lambda$. From Prop. 8, $\Gamma(\nu)$ is nonempty and let $(h,\beta)$ be in $\Gamma(\nu)$. Then, we define for all $s\in\mathbb{R}$ and $t\in\mathbb{R}_{+}$, $\gamma(s,t):=\min\{s,h(\beta^{t})\}$ and for all $x\in\mathbb{R}^{d}$, $\theta(x):=\sup_{k\in\mathbb{N}}\varphi(T^{k}(x))$. It is clear that $\gamma\in\mathcal{KL}_{\rm gen}$ as $\beta\in(0,1)$. Now as $\nu\in\Lambda$, $\nu_{\rm opt}=\sup_{n\in\mathbb{N}}\nu_{n}\in\mathbb{R}$. Moreover, we have $\nu_{\rm opt}=\sup_{n\in\mathbb{N}}\sup_{x\in X^{\mathrm{in}}}\varphi(T^{n}(x))=\sup_{x\in X^{\mathrm{in}}}\sup_{n\in\mathbb{N}}\varphi(T^{n}(x))=\overline{\theta_{X^{\mathrm{in}}}}$. By definition of $\theta$, we have for all $x\in X^{\mathrm{in}}$, for all $k\in\mathbb{N}$, $\varphi(T^{k}(x))\leq\theta(x)$. By definition of $(h,\beta)$, we have for all $k\in\mathbb{N}$, $\nu_{k}\leq h(\beta^{k})$. This is the same as for all $x\in X^{\mathrm{in}}$ and for all $k\in X^{\mathrm{in}}$, $\varphi(T^{k}(x))\leq h(\beta^{k})$. We conclude that $\varphi(T^{k}(x))\leq\min\{\theta(x),h(\beta^{k})\}$ for all $x\in X^{\mathrm{in}}$ and for all $k\in\mathbb{N}$. Note that for a function $\gamma$ constructed as below, for all $s\in\mathbb{R}$, $\inf_{t\geq 0}\gamma(s,t)=\inf_{t\geq 0}\min\{s,h(\beta^{t})\}=\min\{s,h(0)\}$. This follows from the continuity of $h$, the fact $h$ is strictly increasing and $\beta\in(0,1)$.
If moreover, ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$, then from Lemma 10, there exists $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ useful for $\nu$. We could define $\gamma$ using this function $h$. Now, let $k\in\mathcal{S}(\nu,h)$. As $\nu_{k}\leq\nu_{\rm opt}=\overline{\theta_{X^{\mathrm{in}}}}$, we conclude that $\inf_{t\geq 0}\gamma(\overline{\theta_{X^{\mathrm{in}}}},t)=\min\{\overline{\theta_{X^{\mathrm{in}}}},h(0)\}=h(0)$ which ends the proof. ∎
Example 4.
Let us come back to the analysis of the peaks computation problem of Subsection 2.2. We propose to compute a useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound $(\gamma,\theta)$ for the triplet $(X^{\mathrm{in}},T,\varphi)$ defined at Subsection 2.2.
Following the proof of Th 5, one can define the $\mathcal{KL}_{\rm gen}$ function as $(s,t)\mapsto\min\{s,h(\beta^{t})\}$ where $(h,\beta)$ is a pair proposed at Example 2 and $\theta\in\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ as $\mathbb{R}^{2}\ni x\mapsto\sup_{n\in\mathbb{N}}\varphi(T^{n}(x))$.
Another possible pair consists in the $\mathcal{KL}_{\rm gen}$ function: ×$\mathbb{R}\times\mathbb{R}_{+}\ni(s,t)\mapsto\min\{s,(p^{2}(n_{0}+1)(3(t+1))^{-1}\}$ and the constant function in $\operatorname{Fin}\left(X^{\mathrm{in}}\right)$: $\mathbb{R}^{2}\ni x\mapsto p^{2}/3$. We recall that $n_{0}$ is the smallest integer $n$ such that $\varphi_{n}(\alpha)<0$. We can easily check that $\varphi(T^{k}(x))\leq\min\{p^{2}/3,(p^{2}(n_{0}+1)(3(k+1))^{-1}\}$ for all $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$. Moreover, as $\inf_{t\geq 0}\min\{p^{2}/3,(p^{2}(n_{0}+1)(3(t+1))^{-1}\}=0$, there exists $k\in\mathbb{N}$ such that $\nu_{k}>\inf_{t\geq 0}\min\{p^{2}/3,(p^{2}(n_{0}+1)(3(t+1))^{-1}\}$.
In the proof of Th. 5, we exploit Prop. 8 and Prop. 10 which link the nonemptiness of the sets $\mathbb{\Gamma}(u)$, $\mathcal{S}(u,h)$ and ${\mathrm{R}}_{u}^{s}$. We can reformulate Th 5 as Corollary 3.
Corollary 3.
The set $\mathbb{\Gamma}(\nu)$ is nonempty if and only if there exists a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$. If, moreover, there exists $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ useful for $\nu$ then there exists a useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$.
In Th 5 and in Corollary 3, we construct a useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ from a useful $(h,\beta)\in\mathbb{\Gamma}(\nu)$. Now, the question is to know whether is it possible to construct a useful $(h,\beta)\in\mathbb{\Gamma}(\nu)$ from a useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$. This is exactly what was done at Th 4 from the Sontag’s result. Thus, we provide a positive answer. The result is based on an extension of the Sontag’s result for the class $\mathcal{KL}_{\rm gen}$. This permits to fix the value $\beta\in(0,1)$ to only consider problem on functions in $\Omega([0,1])$.
Theorem 6 ($\mathcal{KL}_{\rm gen}$ to $\mathbb{\Gamma}(\nu)$).
Suppose that there exists a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$. Then there exists $h\in\Omega([0,1])$ such that $(h,e^{-1})\in\Gamma(\nu)$. If moreover, the $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ is useful for $\nu$ then $h$ can be chosen to be useful for $\nu$.
Lemma 2.
Let $f:\mathbb{R}_{+}\to\mathbb{R}$ be a decreasing function and let $m\geq\inf_{\mathbb{R}_{+}}f$ ($m$ can be equal to $\infty$- ∞ if $\inf_{\mathbb{R}_{+}}f$ is). There exists a function $g:\mathbb{R}_{+}\mapsto\mathbb{R}$ strictly decreasing and continuous such that:
  1. 1.
    $f\leq g$;
  2. 2.
    $\inf_{\mathbb{R}_{+}}g=m$.
Proof.
If $m\geq f(0)$, the function $x\mapsto g(x)=m+1/(x+1)$ satisfies the above properties. Note that this case includes the case where $f$ is constant on $\mathbb{R}_{+}$.
Now assume that $f(0)>m\geq\inf_{\mathbb{R}_{+}}f$. We denote for all couple of pairs $\{(x,y),(z,t)\}$ where $x,y,z,t\in\mathbb{R}$ by $\mathrm{Aff}\{(x,y),(z,t)\}$ the unique affine function $\ell:\mathbb{R}\mapsto\mathbb{R}$ such that $\ell(x)=y$ and $\ell(z)=t$. We define the functions $h_{0}$ and $h_{n}$ for all $n\in\mathbb{N}^{*}$ as follows :
$F_{0}=f(0)\mathbb{1}_{[0,1]}\text{ and }F_{n}=f(n-1)\mathbb{1}_{\left[n;n+\frac{1}{2}\right]}+\mathrm{Aff}\left\{\left(n+\frac{1}{2};f(n-1)\right);\left(n+1;f(n)\right)\right\}\mathbb{1}_{\left[n+\frac{1}{2};n+1\right]}$
Finally, we define, for all $x\in\mathbb{R}_{+}$, the function $F$:
$F(x)=\sum_{n\in\mathbb{N}}F_{n}(x)+\dfrac{1}{x+1}\text{ and note that}\sum_{n\in\mathbb{N}}F_{n}(x)=\left\{\begin{array}[]{lr}f(0)&\text{ if }x\in[0,1]\\ f(\lfloor x\rfloor-1)&\text{ if }x\in[\lfloor x\rfloor,\lfloor x\rfloor+1/2]\\ \ell\left(x\right)&\text{ if }x\in(\lfloor x\rfloor+1/2,\lfloor x\rfloor+1)\end{array}\right.$
where $\ell=\mathrm{Aff}\left\{\left(\lfloor x\rfloor+1/2;f(\lfloor x\rfloor-1)\right);\left(\lfloor x\rfloor+1;f(\lfloor x\rfloor)\right)\right\}$.
The function $\sum_{n\in\mathbb{N}}F_{n}$ is clearly continuous and decreasing (as $f$ does) on $\mathbb{R}_{+}$. The function $\sum_{n\in\mathbb{N}}F_{n}$ is not strictly decreasing but adding $x\mapsto(x+1)^{-1}$ makes $F$ to be strictly decreasing and continuous. We also have $f\leq\sum_{n\in\mathbb{N}}F_{n}$ as $f$ is decreasing. As $x\mapsto(x+1)^{-1}$ is strictly positive we have $f\leq F$. Finally, we have $\inf_{\mathbb{R}_{+}}F=\inf_{\mathbb{R}_{+}}f$. Indeed, for all $n\in\mathbb{N}^{*}$ and all $x\in[n,n+1)$, $F_{n}\leq f(n-1)$, it implies that $\inf_{\mathbb{R}_{+}}F=\inf_{x\in\mathbb{R}_{+}}\inf_{n\in\mathbb{N}^{*}}F_{n}(x)\leq\inf_{n\in\mathbb{N}^{*}}f(n-1)=\inf_{\mathbb{R}_{+}}f$ (as $f$ decreases). The reverse inequality holds as $f\leq F$.
The function $F$ does not necessarily satisfy $\inf_{\mathbb{R}_{+}}h=m$. Hence, we define for all $x\in\mathbb{R}_{+}$:
$g(x)=\left\{\begin{array}[]{lr}\max\{F(x),(x+1)^{-1}+m\}&\text{ if }m>\inf_{\mathbb{R}_{+}}f\\ F(x)&\text{ if }m=\inf_{\mathbb{R}_{+}}f\end{array}\right.$
Suppose that $m>\inf_{\mathbb{R}_{+}}f$. The function $g$ is clearly strictly decreasing and continuous and satisfies $f\leq g$ as $F$ does. As $m>\inf_{\mathbb{R}_{+}}f=\inf_{\mathbb{R}_{+}}F$, there exists $x_{0}$ such that $F(x)<m$ for all $x\geq x_{0}$ and then $g(x)=(x+1)^{-1}+m$ for all $x\geq x_{0}$. This implies that $\inf_{\mathbb{R}_{+}}g=m$. ∎
Lemma 3 (Sontag’s extension).
Let $\gamma\in\mathcal{KL}_{\rm gen}$. Let $s\in\mathbb{R}$ and let $m\geq\inf_{t\geq 0}\gamma(s,t)$ be finite. Then there exists $h\in\Omega([0,1])$ such that for all $t\geq 0$, $\gamma(s,t)\leq h(e^{-t})$ and $h(0)=m$.
Proof.
By definition, $t\mapsto\gamma(s,t)$ is decreasing and applying Lemma 2, there exists a function $\sigma:\mathbb{R}_{+}\mapsto\mathbb{R}$ strictly decreasing and continuous such that $\gamma(s,t)\leq\sigma(t)$ for all $t\geq 0$ and $\inf_{\mathbb{R}_{+}}\sigma=m$. Let us define for all $r\in(0,1]$, $h(r)=\sigma(-\ln r)$ and $\lim_{x\to 0^{+}}h(x)=\lim_{y\to+\infty}\sigma(y)=m$. Then defining $h(0)=m$ makes $h$ continuous at 0. Finally, $h$ belongs to $\Omega([0,1])$. ∎
Proof of Th. 6.
By assumption, there exist a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound $(\gamma,\theta)$ for $(X^{\mathrm{in}},T,\varphi)$. Let $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$. This implies that $\varphi(T^{k}(x))\leq\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)$. By Lemma 3, there exists $h\in\Omega([0,1])$ such that $\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)\leq h(e^{-k})$ for all $k\in\mathbb{N}$. This leads to $(h,e^{-1})\in\mathbb{\Gamma}(\nu)$. If there exists $k\in\mathbb{N}$ such that $\inf_{t\in\mathbb{R}}\gamma(\overline{\theta_{X^{\mathrm{in}}}},t)<\nu_{k}$, we can choose, as it is stated in Lemma 3, $h$ such that $h(0)=m\in(\inf_{t\in\mathbb{R}}\gamma(\overline{\psi},t),\nu_{k})$. ∎
We extend the class of $\mathcal{K}\mathcal{L}$ functions which classically serve for dynamical systems stability. In our context, we have a similar use and our $\mathcal{KL}_{\rm gen}$ class constitutes the main part of $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bounds of $\varphi(T^{k}(x))$. We prove that such upper bounds exist if and only if $\nu\in\Lambda$. Moreover, if ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$, we are able to construct a function $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ useful for $\nu$ and such that $e^{-1}\in\mathbb{\Gamma}_{\rm sc}(\nu,h)$.
Unfortunately, in practice, to prove that a pair $(\gamma,\theta)$ is a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ seems to be as difficult as proving that a pair $(h,\beta)$ belongs to $\mathbb{\Gamma}(\nu)$. We thus propose a new type of Lyapunov functions for which the conditions does not involve the whole trajectories of the analyzed dynamical system.
5 Lyapunov type Functions for Discrete-time Systems Maximization
Classically, Lyapunov functions are used as certificates of stability (see  [Ela05] or [KP01]). They bring a constructive approach to prove the stability of the dynamical system through converse Lyapunov theorems which prove that stability notions are equivalent to the existence of Lyapunov functions [Kha02]. For some linear, polynomial or piecewise polynomial systems, Lyapunov functions can be computed from numerical optimization solvers based on linear or semidefinite programming (for example see [GH15] and references therein).
Classical quadratic Lyapunov functions solve our problem when the dynamical system is stable and affine and the objective function is quadratic [Adj21]. In Subsection 5.1, we generalize this approach theoretically. Then in Subsection 5.2, we define a more general class of Lyapunov functions. The existence of those functions does not ensure the stability of the system but just the finiteness of the limsup of the sequence $\nu$ defined at Eq. (1). Last but not least, we etablish an equivalence theorem between the existence of those new Lyapunov functions and the existence of an element $(h,\beta)\in\mathbb{\Gamma}(\nu)$.
5.1 Classical Lyapunov Functions Theory
5.1.1 Definition and Useful Facts
[…]
Definition 7 (Classical Lyapunov Functions).
A function $V:\mathbb{R}^{d}\to\mathbb{R}_{+}$ is said to a Lyapunov function if and only if :
  1. 1.
    There exist $\alpha_{1},\alpha_{2}\in\mathcal{K}_{\infty}$ such that:
    $\alpha_{1}\circ\|\cdot\|\leq V\leq\alpha_{2}\circ\|\cdot\|$
  2. 2.
    There exists $\lambda\in(0,1)$ such that
    $V\circ T\leq\lambda V$
[…]
  1. 1.
    there exists $\alpha:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ positive definite111a function $g:X\mapsto\mathbb{R}_{+}$ where $X\subset\mathbb{R}^{d}$ and $0\in X$ is said to be positive definite (on $X$) if $g(0)=0$ and $g(x)>0$ for all $x\in X\backslash\{0\}$ and continuous such that $V\circ T\leq V-\alpha\circ\|\cdot\|$;
  2. 2.
    there exists $\rho:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ positive definite, continuous and such that $\mathrm{Id}-\rho$ is positive definite satisfying $V\circ T\leq\rho\circ V$.
We will propose a proof for this equivalence later on the new class of Lyapunov functions used for the maximization problem.
First, we remark that from the first statement of Def. 7, $V$ is positive definite. If we require $V$ to be continuous and such that $\lim V(x)=+\infty$ as $\|x\|\to+\infty$ (see e.g. [Adj24]), then the first statement of Def. 7 is actually equivalent to $V$ is postive definite. Second, if there exists a classical Lyapunov function for $T$, we $T$ must be null at zero [^2] Actually, if $T(e)=e$ for a nonzero vector, we could replace $V$ by $V(\cdot-e)$ and the lower and the upper bounds on $V(\cdot-e)$ by respectively $\alpha_{1}(\|\cdot-e\|)$ and $\alpha_{2}(\|\cdot-e\|)$. Finally, from the second statement of Def 7, we can define a kind of norm operator relative to a Lyapunov function. We have, for all Lyapunov function $V$ associated to the dynamic $T$, for all $k\in\mathbb{N}$:
$N_{T^{k}}(V):=\sup_{\begin{subarray}{c}x\in\mathbb{R}^{d}\\ x\neq 0\end{subarray}}\dfrac{V(T^{k}(x))}{V(x)}\in(0,1)$ (5)
We will simply write $N_{T}(V)$ when $k=1$. By definition, we have for all $x\in\mathbb{R}^{d}$, $V(T(x))\leq N_{T}(V)V(x)$. Using this and the recurrence principle, we get, for all $k\in\mathbb{N}$, $N_{T^{k}}(V)\leq(N_{T}(V))^{k}$.
5.1.2 A Direct Classical Lyapunov Theorem
Let $f$ be a function from $\mathbb{R}^{d}$ to $\mathbb{R}$. We recall that the function from $\mathbb{R}_{+}$ to $\mathbb{R}$:
$\alpha^{f}:s\mapsto\sup_{\|y\|\leq s}f(y)$ (6)
is continuous when $f$ is continuous (see e.g. [Kha02] or extensions in [Adj24]). Obviously, the function $\alpha^{f}$ is increasing. Moreover, if $f(0)=0$ then $\alpha^{f}$ is nonnegative and satisfies $\alpha^{f}(0)=0$. Last but not least, we have for all $x\in\mathbb{R}^{d}$, $f(x)\leq\alpha^{f}(\|x\|)$.
Theorem 7 (Classical Lyapunov to $\mathbb{\Gamma}(\nu)$).
Assume that $\varphi$ is continuous and there exists a Lyapunov function for $T$. Then there exists $(h,\beta)\in\mathbb{\Gamma}(\nu)$ such that $h(0)=0$. Thus, if Assumption 2 holds, $h$ is useful for $\nu$.
Proof.
If there exists a Lyapunov function $V$ for $T$, there exists $\alpha_{1}\in\mathcal{K}_{\infty}$ such that $\alpha_{1}\circ\|\cdot\|\leq V$. Recall that we have assumed that $\varphi(0)=0$. With the continuity assumption of $\varphi$, we conclude that $\alpha^{\varphi}$ (defined at Eq. (6)) is nonnegative, increasing, continuous and is equal to 0 at 0. Two small issues : $\alpha^{\varphi}$ is not guaranteed to be strictly increasing and surjective from $\mathbb{R}_{+}$ to $\mathbb{R}_{+}$ (the limit of $\alpha^{\varphi}$ at +$\infty$+ ∞ can be finite). We thus add to $\alpha^{\varphi}$ a function $\psi\in\mathcal{K}_{\infty}$ and we have for $y\in\mathbb{R}^{d}$, $\varphi(y)\leq(\alpha^{\varphi}+\psi)(\|y\|)$ and $\alpha:=\alpha^{\varphi}+\psi$ belongs to $\mathcal{K}_{\infty}$.
From $\alpha_{1}\circ\|\cdot\|\leq V$ and $\alpha_{1}\in\mathcal{K}_{\infty}$, we get $\varphi\leq\alpha\circ\displaystyle{{\alpha_{1}}_{\left(\mathbb{R}_{+}\right)}^{-1}}\circ V$ as $\alpha$ is increasing. Moreover, we have, for all $x\in X^{\mathrm{in}}$, for all $k\in\mathbb{N}$, $V(T^{k}(x))\leq N_{T^{k}}(V)V(x)\leq(N_{T}(V))^{k}V(x)\leq(N_{T}(V))^{k}\overline{V_{X^{\mathrm{in}}}}$. As $\alpha$ and ${{\alpha_{1}}_{\left(\mathbb{R}_{+}\right)}^{-1}}$ are increasing, we have for all $k\in\mathbb{N}$, $\nu_{k}\leq\alpha\left(\displaystyle{{\alpha_{1}}_{\left(\mathbb{R}_{+}\right)}^{-1}}\left(N_{T}(V))^{k}\overline{V_{X^{\mathrm{in}}}}\right)\right)$. We thus define $h:x\mapsto\alpha\left(\displaystyle{{\alpha_{1}}_{\left(\mathbb{R}_{+}\right)}^{-1}}(x\overline{V_{X^{\mathrm{in}}}})\right)$ and $\beta=N_{T}(V)$. We have $h(0)=0$ and we conclude that $(h,\beta)\in\mathbb{\Gamma}(\nu)$ is useful for $\nu$ when Assumption 2 holds. ∎
The goal of extending the concept of Lyapunov functions is to weaken the hypothesis of Th 7 by removing the continuity and the positivity assumptions on $\varphi$. The extension also allows to prove the equivalence between the existence of useful pair $(h,\beta)\in\mathbb{\Gamma}(\nu)$ and the existence of this new type of Lyapunov functions.
5.2 Opt-Lyapunov Functions for Discrete-time Systems Maximization
5.2.1 Definitions and Useful Facts
We recall that the domain of function $g:\mathbb{R}^{d}\mapsto(-\infty,+\infty]$ is the set of $x\in\mathbb{R}^{d}$ such that $f(x)<+\infty$. We denote by $\operatorname{dom}(f)$ the domain of a extended real valued function.
Definition 8 (Opt-Lyapunov functions).
A function $V$ is said to be an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ if and only if:
  1. 1.
    $V:\mathbb{R}^{d}\mapsto[0,+\infty]$;
  2. 2.
    $\overline{V_{X^{\mathrm{in}}}}\in(0,1]$;
  3. 3.
    $V\circ T\leq\lambda V$ for some $\lambda\in(0,1)$.
Example 5.
Let us consider the dynamical system of the example of Subsection 2.2. Recall that $\mu\in(0,1/3)$ and the set of initial conditions is :
$X^{\mathrm{in}}=([2\mu,1]\times[\mu,1])\cap\operatorname{Vect}\{(2,1)^{\intercal}\}$
and the dynamics is the following:
$\mathbb{R}^{2}\ni(x,y)\mapsto T(x,y)=A\begin{pmatrix}x\\ y\end{pmatrix}\text{ where }A=\begin{pmatrix}1&1\\ \frac{1}{4}&1\end{pmatrix}$
The function defined, for all $x=(x_{1},x_{2})^{\intercal}\in\mathbb{R}^{2}$ as follows:
$V(x)=\left\{\begin{array}[]{cr}\dfrac{2\mu^{2}}{x_{1}x_{2}}&\text{ if }x\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}\\ \\ 0&\text{ otherwise}\end{array}\right.$
is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
Let us prove that it is true. Let $x\in\mathbb{R}^{2}$. If $x\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}$, then $x=(2,1)^{\intercal}y$ with $y\in\mathbb{R}^{*}$. First, the coordinates of $x$ has the same sign and so $V$ is nonnegative. The fact that $V(x)$ is nonnegative when for $x\notin\operatorname{Vect}\{(2,1)^{\intercal}\}$ or $x=0$ is obvious. Second, let $x\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}$ and let $y\neq 0$ such that $x=(2,1)^{\intercal}y$. Then, $V(T(x))=V((3y,3y/2)^{\intercal})=(4\mu^{2})/(9y^{2})$ and $V(x)=\mu^{2}/y^{2}$ then $V(T(x))/V(x)=(4y^{2})/(9y^{2})=4/9$. Any value in $[4/9,1)$ can be chosen. Let us take $5/9$. As we have $T(z)\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}$ if and only if $z\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}$, we conclude that for all $z\notin\operatorname{Vect}\{(2,1)^{\intercal}\}$ or $z=0$, we have $V(T(z))=V(z)=0$. We conclude that $V(T(z))\leq 5V(z)/9$ for all $z\in\mathbb{R}^{2}$. Finally, since $\mu>0$, we have $\max_{x\in X^{\mathrm{in}}}V(x)=\max_{y\in[\mu,1/2]}(2\mu^{2})/(2y^{2})^{-1}=(2\mu^{2})(2\mu^{2})^{-1}=1$.
Proposition 14.
If $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ then :
  1. 1.
    for all $k\in\mathbb{N}$, $T^{k}(X^{\mathrm{in}})\subseteq V^{-1}([0,1])\subseteq\operatorname{dom}(V)$;
  2. 2.
    for all fixed points $x$ of $T$, $V(x)\in\{0,+\infty\}$.
Proof.
1. From the second statement for all $x\in X^{\mathrm{in}}$, $V(x)\in[0,1]$. From the third statement and the fact that $V$ is nonnegative, we have, for all $x\in X^{\mathrm{in}}$, for all $k\in\mathbb{N}$, $0\leq V(T^{k}(x))\leq V(x)\leq 1$.
2. Let $x$ be a fixed point of $T$. If $V(x)=+\infty$, there is nothing to prove. Now, suppose that $V(x)$ is finite. From the third statement, $0\leq V(T(x))=V(x)\leq\lambda V(x)$ for some $\lambda\in(0,1)$. Hence, $V(x)(1-\lambda)=0$ and we conclude that $V(x)=0$. ∎
Corollary 4.
If $X^{\mathrm{in}}$ is included in the fixed point set of $T$ then a function $V:\mathbb{R}^{d}\mapsto[0,+\infty]$ such that $V\circ T\leq\lambda V$ for all some $\lambda\in(0,1)$ cannot satisfy $\overline{V_{X^{\mathrm{in}}}}>0$.
Corollary 4 extends the fact that if a classical Lyapunov function (as defined in Def 7) exists for a self-map $T$, $T$ cannot have non-zero fixed point. If an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ exists, $X^{\mathrm{in}}$ contains a vector $x$ such that $x\neq f(x)$.
Prop. 15 shows that the set of Opt-Lyapunov functions contains the classical Lyapunov functions.
Proposition 15 (Classical is Opt).
If $V$ is a classical Lyapunov function for $T$ and $X^{\mathrm{in}}$ is not reduced to $\{0\}$, then $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
Actually, we can replace equivalently the second statement of Def. 8 by $\overline{V_{X^{\mathrm{in}}}}$ is finite and strictly positive. Indeed, with this condition, we recover an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ with $\overline{W_{X^{\mathrm{in}}}}\in(0,1]$ by taking $W:=V/\overline{V}$.
Proposition 16.
The existence of a function $W:\mathbb{R}^{d}\mapsto[0,+\infty]$ such that :
  1. 1.
    $\overline{W_{X^{\mathrm{in}}}}$ is strictly positive and finite ;
  2. 2.
    $W\circ T\leq\rho\circ W$ for a function $\rho:[0,+\infty]\to[0,+\infty]$ such that:
    • $\rho(x)=0\iff x=0$ and $\rho(x)=+\infty\iff x=+\infty$;
    • $\rho_{\mid\mathbb{R}_{+}}\in\mathcal{K}$;
    • $\rho(x)<x$ for all $x\in\mathbb{R}_{+}^{*}$.
is equivalent to the existence of a Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
Lemma 4.
Let $\gamma$ be strictly greater than 1. Let $f\in\mathcal{K}_{\infty}$ verifying for all $x>0$, $f(x)>x$. Then, there exists a function $g\in\mathcal{K}$ which satisfies $g(f(x))=\gamma g(x)$ for all $x\in\mathbb{R}_{+}$.
Lemma 4 is a slighty modified version of [Kel14, Lemma 25] for which $f-\mathrm{Id}$ is supposed to be in $\mathcal{K}_{\infty}$. To be self-contained, we propose in Appendix a proof inspired by the one provided in [Kel14].
Lemma 5.
Let $\lambda\in(0,1)$. Let $f\in\mathcal{K}$ verifying for all $x>0$, $f(x)<x$. Then, there exists a function $g\in\mathcal{K}$ satisfying $g(f(x))=\lambda g(x)$ for all $x\in\mathbb{R}_{+}$.
Again a proof of Lemma 5 is given in [Kel14]. A proof is provided in Appendix for the convenience of the reader.
Proof of Prop. 16.
Suppose that such $W$ and $\rho$ exist. Let $\lambda\in(0,1)$. From Lemma 5, there exists $g$ such that $g(\rho(x))=\lambda g(x)$ for all $x\in\mathbb{R}_{+}$. By extension, we define $g(+\infty)=+\infty$. Then as $g$ is increasing, we have $g(W(T(x)))\leq g(\rho(W(x))=\lambda g(W(x))$ for all $x\in\mathbb{R}^{d}$. Now, as $g$ is strictly increasing there exists some $x\in X^{\mathrm{in}}$ such that $g(W(x))$ is strictly positive and $\overline{g\circ W_{X^{\mathrm{in}}}}$ is strictly positive and finite. Finally, $(g\circ W)/\overline{g\circ W_{X^{\mathrm{in}}}}$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
The reverse implication is obvious with $\rho(x)=\lambda x$ for all $x\in\mathbb{R}_{+}$. ∎
5.2.2 Operator Ratios From Proper Positive Semidefinite Functions
We extend the concept of operator ratios (defined at Eq. (5)) for which the arguments are positive semidefinite functions rather than positive definite functions. The arguments can also be functions which take infinite values. So, let us denote by $\mathcal{PSD}$ the set of nonzero proper positive semidefinite functions from $\mathbb{R}^{d}$ to $[0,+\infty]$ i.e. :
$\mathcal{PSD}:=\{f:\mathbb{R}^{d}\mapsto[0,+\infty]:\exists\,x\in\mathbb{R}^{d},\ f(x)\in\mathbb{R}_{+}^{*}\}$ (7)
We are able to extend the operator ratio function to $\mathcal{PSD}$ functions.
Definition 9 (Operator Ratios PSD Based).
Let $U:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$ and $k\in\mathbb{N}$. We associate to $U$ and $k$ the operator ratio function $\overline{N_{U^{k}}}:\mathcal{PSD}\mapsto[0,+\infty]$ defined as follows:
$\overline{N_{U^{k}}}(P):=\sup_{P(x)\in\mathbb{R}_{+}^{*}}\dfrac{P(U^{k}(x))}{P(x)}$
As for $N_{U^{k}}$, we will simply write $\overline{N_{U}}$ for $k=1$.
Recall that $\mathrm{Id}$ denotes the identity function $\mathbb{R}^{d}\ni x:\mapsto x$. It is straightforward to see that, if $P\in\mathcal{PSD}$, $\overline{N_{\mathrm{Id}}}(P)=1$.
Example 6.
We proved at Example 5 that:
$\overline{N_{T}}(V)=\sup_{\{x\in\mathbb{R}^{2}:V(x)>0\}}\dfrac{V(T(x))}{V(x)}=\max_{x\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}}\dfrac{V(T(x))}{V(x)}=4/9.$
The operator function is well-defined but the inequality $\overline{N_{U^{k}}}(P)\leq\overline{N_{U}}(P)^{k}$ can fail to be true as we can see at Example 7.
Example 7 (A counterexample to the power commutation).
Let us consider the function $U:x\mapsto-x$ and $P:x\mapsto\max\{x,0\}$. We have $\overline{N_{U}}(P)=\sup_{P(x)>0}P(U(x))/P(x)=\sup_{x>0}P(-x)/x=\sup_{x>0}0/x=0$ whereas $\overline{N_{U^{2}}}(P)=\sup_{x>0}P(x)/x=\sup_{x>0}x/x=1$. Finally, we have $\overline{N_{U^{2}}}(P)>\overline{N_{U}}(P)^{2}$.
To obtain the inequality $\overline{N_{U^{k}}}(P)\leq\overline{N_{U}}(P)^{k}$, we have to consider a smaller class of operators $P$ and we introduce for a self-map $U:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$, the set $\mathcal{D}\left(U\right)$ of operators defined as follows:
$\mathcal{D}\left(U\right):=\{P\in\mathcal{PSD}:\forall\,x\in\mathbb{R}^{d},\ P(U(x))=0\implies P(U^{2}(x))=0\}$ (8)
We remark that $P\in\mathcal{D}\left(U\right)$ if and only if $P(U^{2}(x))\in(0,+\infty]\implies P(U(x))\in(0,+\infty]$. Moreover, if $P\in\mathcal{D}\left(U\right)$, this leads to for all $k\geq 1$, $P(U^{k+1}(x))\in(0,+\infty]\implies P(U^{k}(x))\in(0,+\infty]$ and we also have if $P(U^{k}(x))\in(0,+\infty]$ for some $k\geq 1$ then $P(U^{j}(x))\in(0,+\infty]$ for all $1\leq j\leq k$.
Example 8 ($P$ not in $\mathcal{D}\left(U\right)$).
Let consider $U(x)=-x$ and $P(x)=\max\{x,0\}$. Let $x$ be a strictly positive real. Then, $P(U(x))=P(-x)=0$. However, $P(U^{2}(x))=P(x)=x>0$. So $P\notin\mathcal{D}\left(U\right)$.
Proposition 17.
Let $U:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$ and $P\in\mathcal{D}\left(U\right)$. Then, for all $k\in\mathbb{N}^{*}$, $\overline{N_{U^{k}}}(P)\leq\left(\overline{N_{U}}(P)\right)^{k}$ and if $\overline{N_{U}}(P)$ is finite this inequality holds for $k=0$.
Proof.
For $k=0$, the result holds as $\overline{N_{\mathrm{Id}}}(P)=1$ if there exists $x\in\mathbb{R}^{d}$ such that $P(x)>0$ and 0 otherwise. If $\overline{N_{U}}(P)$ is finite $(\overline{N_{U}}(P))^{0}$ is defined and equal to 1.
Now let $k\in\mathbb{N}^{*}$. If $\overline{N_{U}}(P)=+\infty$, the inequality is obviously true. Suppose now that $\overline{N_{U}}(P)<+\infty$. From the defintion of $\overline{N_{U}}(P)$, we have for all $x\in\mathbb{R}^{d}$ and all $j\in\mathbb{N}^{*}$ such that $P(U^{j-1}(x))\in\mathbb{R}_{+}^{*}$, $P(U^{j}(x))\leq\overline{N_{U}}(P)P(U^{j-1}(x))$. From the fact that $P\in\mathcal{D}\left(U\right)$, we also have for all $x\in\mathbb{R}^{d}$ and all $j\in\mathbb{N}^{*}$ such that $P(U^{j}(x))=0$, $P(U^{j+1}(x))\leq\overline{N_{U}}(P)P(U^{j}(x))$ as it implies that $P(U^{j+1}(x))=0$. Finally, we have, for all $x\in\mathbb{R}^{d}$ such that $P(x)>0$, $P(U^{k}(x))\leq\overline{N_{U}}(P)P(U^{k-1}(x))$. By dividing by $P(x)$ and taking the supremum, we obtain $\overline{N_{U^{k}}}(P)\leq\overline{N_{U}}(P)\overline{N_{U^{k-1}}}(P)$ and we conclude by recurrence. ∎
Proposition 18.
If $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$, then $V\in\mathcal{D}\left(U\right)$ and $\overline{N_{T}}(V)\in[0,1)$.
Proof.
First $V^{-1}(\mathbb{R}_{+}^{*})$ is nonempty from the second statement of Def. 8. The third statement implies that $V(T^{2}(x))\leq\lambda V(T(x))$ for all $x\in\mathbb{R}^{d}$ for some $\lambda\in(0,1)$ and thus $V(T^{2}(x))>0$ implies that $V(T(x))>0$. We conclude that $V\in\mathcal{D}\left(T\right)$. The third statement also implies that $V(T(x))\leq\lambda V(x)$ for all $x\in\operatorname{dom}(V)$ for some $\lambda\in(0,1)$ and thus $N_{T}(V)\leq\lambda<1$. It can be equal to 0 if and only if for all $x$ such that $V(x)>0$, $V(T(x))=0$. ∎
To prove that a function $V\in\mathcal{PSD}$ is an Opt-Lyapunov for $X^{\mathrm{in}},T)$, we could use any value greater $\overline{N_{T}}(V)$. The computation of $\overline{N_{T}}(V)$ ensures to take the best value possible. It is important in the context of using the function $\mathfrak{F}_{u}(k,h,\beta)$ with an Opt-Lyapunov function since $\beta\mapsto\mathfrak{F}_{u}(k,h,\beta)$ is decreasing (Prop. 12).
Proposition 19.
Let $U:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$. Suppose that there exists a classical Lyapunov function $V$ for $U$. Then, $V\in\mathcal{D}\left(U\right)$ and $\overline{N_{U}}(V)=N_{U}(V)$.
Proof.
If there exists a Lyapunov function for $U$, then $U(0)=0$. From the double inequalities $\alpha\circ\|\cdot\|\leq V\leq\beta\circ\|\cdot\|$ and thus $V(x)=0$ iff $x=0$. Thus if $V(U^{2}(x))>0$ then $U^{2}(x)\neq 0$ and thus $U(x)\neq 0$ hence $V(U(x))>0$ and $V\in\mathcal{D}\left(U\right)$. The equality between $\overline{N_{U}}(V)$ and $N_{U}(V)$ follows readily from the fact that $\{y\in\mathbb{R}^{d}:V(y)>0\}=\mathbb{R}^{d}\backslash\{0\}$. ∎
5.2.3 $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility and $(X^{\mathrm{in}},T,\varphi)$-compatibility
In the proof of Th. 7, the construction of $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ from a classical Lyapunov function $V$ is based on the existence of a function $\alpha\in\mathcal{K}_{\infty}$ such that $\varphi\leq\alpha\circ V$. The existence of this latter comes from the fact that in Th. 7, $\varphi$ is continuous and, by definition of $V$, $\alpha^{\prime}\circ\|\cdot\|\leq V$ for some $\alpha^{\prime}\in\mathcal{K}_{\infty}$. However, when $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$, $V$ is not necessarily positive definite and we cannot guarantee the existence of $\alpha\in\mathcal{K}_{\infty}$ such that $\alpha\circ\varphi\leq V$. We have to use a different class of function to find a similar relation between $\varphi$ and $V$. First, we propose to restrict the interval where a function $\alpha$ should be continuous and strictly increasing. Indeed, the analysis only concerns the values $\{\varphi(T^{k}(x)),k\in\mathbb{N},x\in X^{\mathrm{in}}\}$ and $\{V(x),x\in X^{\mathrm{in}}\}$. Hence, we introduce the following interval :
$I_{X^{\mathrm{in}},T}^{\varphi}:=\overline{\operatorname{conv}}(\{\varphi(T^{k}(x)):k\in\mathbb{N},\ x\in X^{\mathrm{in}}\})\in\mathbb{I}\left(\mathbb{R}\right)$
Proposition 20.
The following assertions hold:
  1. 1.
    For all $k\in\mathbb{N}$, $\nu_{k}\in I_{X^{\mathrm{in}},T}^{\varphi}$. The values, if they are finite, $\limsup_{n\in\mathbb{N}}\nu_{n}$ and $\liminf_{n\in\mathbb{N}}\nu_{n}$ also belong to $I_{X^{\mathrm{in}},T}^{\varphi}$.
  2. 2.
    $I_{X^{\mathrm{in}},T}^{\varphi}=\left[\inf_{k\in\mathbb{N}}\nu_{k},\sup_{k\in\mathbb{N}}\nu_{k}\right]\cap\mathbb{R}$
Proof.
Let $k\in\mathbb{N}$. There exists for all $n\in\mathbb{N}$, $x_{n}\in X^{\mathrm{in}}$ such that $\nu_{k}-1/n\leq\varphi(T^{k}(x_{n}))\leq\nu_{k}$ and thus the sequence $(\varphi(T^{k}(x_{n})))_{n\in\mathbb{N}}$ converges to $\nu_{k}$ and thus $\nu_{k}\in I_{X^{\mathrm{in}},T}^{\varphi}$. The values, if they are finite, $\limsup_{k\in\mathbb{N}}\nu_{k}$ and $\liminf_{k\in\mathbb{N}}\nu_{k}$ are limits of subsequences of $(\nu_{k})_{k\in\mathbb{N}}$ and they belongs to $I_{X^{\mathrm{in}},T}^{\varphi}$. From Prop. 1, if $\limsup_{k\in\mathbb{N}}\nu_{k}$ (resp. $\liminf_{k\in\mathbb{N}}\nu_{k}$) is infinite, the value $\sup_{k\in\mathbb{N}}\nu_{k}$ (resp. $\inf_{k\in\mathbb{N}}\nu_{k}$ is equal to +$\infty$+ ∞ (resp. $\infty$- ∞).
Now, if one of the value $\inf_{k\in\mathbb{N}}\nu_{k}$ or $\sup_{k\in\mathbb{N}}\nu_{k}$ is not finite, the infinity passes to the related bound of $I_{X^{\mathrm{in}},T}^{\varphi}$. Now if those values are both finite, we can extract sequences in $I_{X^{\mathrm{in}},T}^{\varphi}$ which converge respectively to $\inf_{m\in\mathbb{N}}\nu_{m}$ and $\sup_{m\in\mathbb{N}}\nu_{m}$. For example, we can define $\varphi(T^{k_{n}}(x_{k_{n}}))\subset I_{X^{\mathrm{in}},T}^{\varphi}$ such that for all $n\in\mathbb{N}$, $\inf_{m\in\mathbb{N}}\nu_{m}-1/n\leq\nu_{k_{n}}-1/n\leq\varphi(T^{k_{n}}(x_{k_{n}}))\leq\nu_{k_{n}}\leq\inf_{m\in\mathbb{N}}\nu_{m}+1/n$. This sequence converge to $\inf_{m\in\mathbb{N}}\nu_{m}$. ∎
Let $V$ be an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$. To pass from $\alpha\circ\varphi\leq V$ to $\varphi\leq\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\alpha\circ V$, $\alpha$ has to be surjective and strictly increasing as a function of $J\to\{V(x),x\in X^{\mathrm{in}}\}\subseteq[0,1]$. From the earlier discussion, $J$ must also contain $I_{X^{\mathrm{in}},T}^{\varphi}$. Therefore, we define the following set of functions:
$\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}:=\{\alpha:\mathbb{R}\mapsto\mathbb{R}:\exists\,I\in\mathbb{I}\left(\mathbb{R}\right)\text{ s.t. }\alpha(I)=[0,1]\text{ and}\operatorname{conv}(I_{X^{\mathrm{in}},T}^{\varphi}\cup I)\in\mathrm{SIC}(\alpha)\}$
Proposition 21.
Let $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$. There is only one interval $I$ such that $\alpha(I)=[0,1]$ and $I\in\mathrm{SIC}(\alpha)$.
Proof.
Let us consider $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ and $I$ such that $\alpha(I)=[0,1]$. As $\alpha$ is continuous and strictly increasing on $I$, ${{\alpha}_{\left(I\right)}^{-1}}$ exists and is also continuous and strictly increasing from $[0,1]\to I$. Hence, $I$ is compact. Moreover, such interval $I$ is unique. Otherwise, consider $I,J\in\mathrm{SIC}(\alpha)$ for which we have $I\neq J$ and $\alpha(I)=\alpha(J)=[0,1]$. Suppose there exists $x\in I$ such that $x\leq\min J$. We have $\alpha(x)<\alpha(\min J)=0$ but $x\in I$ and $\alpha(x)\geq\alpha(\min I)=0$. ∎
We present some useful properties of functions belonging to $\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ and their relation with $\nu$.
Proposition 22.
For all $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ such that there exists an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$, $V$, satisfying $\alpha(\varphi(T^{k}(x)))\leq V(T^{k}(x))$ for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$, we have $\alpha(\limsup_{k\to+\infty}\nu_{k})\leq 0$.
Proof.
Let $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$. By assumption, $\alpha(\varphi(T^{k}(x)))\leq V(\varphi(T^{k}(x)))\leq\lambda^{k}V(x)$. Taking the sup over $X^{\mathrm{in}}$ and from the increasing condition and the continuity of $\alpha$, $\alpha(\nu_{k})\leq\lambda^{k}\overline{V_{X^{\mathrm{in}}}}$. Taking the limsup, we get $\limsup_{k\mapsto+\infty}\alpha(\nu_{k})=\alpha(\limsup_{k\mapsto+\infty}\nu_{k})\leq 0$. ∎
Corollary 5.
Suppose there exist $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ and an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ $V$ such that $\alpha(\varphi(T^{k}(x)))\leq V(T^{k}(x))$ for all $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$. Then:
  1. 1.
    $\nu\in\Lambda$;
  2. 2.
    If ${\mathrm{R}}_{\nu}^{s}=\emptyset$, for all $\alpha(\varphi(T^{k}(x)))\leq 0$ for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$.
Proof.
The first statement follows readily from Prop. 22. The second statement follows from Prop. 3 and Prop. 22, since we have, if ${\mathrm{R}}_{\nu}^{s}=\emptyset$, for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$, $\varphi(T^{k}(x))\leq\sup_{k\in\mathbb{N}}\nu_{k}=\limsup_{n\to+\infty}\nu_{n}$. ∎
The set $\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ defines the class of functions that we must consider to obtain relations $\alpha\circ\varphi\leq V$ when $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$. Another important consideration to take into account is usefulness of functions $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ based on $\alpha$. An Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ is always nonnegative and if $\alpha(\nu_{k})\leq 0$ for all $k\in\mathbb{N}$, the inequality is obvious. Then $\alpha(\nu_{k})$ must be strictly positive for some $k\in\mathbb{N}$. By Prop. 22 and Corollary 5, we have automatically $k\in{\mathrm{R}}_{\nu}^{s}$. We propose a special definition for those particular functions $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$.
Definition 10 ($(X^{\mathrm{in}},T,\varphi)$-Certificate of compatibility).
Let $g\in\mathcal{PSD}$ such that $\{T^{k}(x):x\in X^{\mathrm{in}},k\in\mathbb{N}\}\subseteq\operatorname{dom}(g)$. A function $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ such that $\alpha(\nu_{k})>0$ for some $k\in{\mathrm{R}}_{\nu}^{s}$ and $\alpha(\varphi(T^{k}(x)))\leq g(T^{k}(x))$ for all $k\in\mathbb{N}$ and all $x\in X^{\mathrm{in}}$ is called a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $g$.
The problem is to decide which conditions on $X^{\mathrm{in}}$, $T$, $\varphi$ and $g\in\mathcal{PSD}$ such that $\{T^{k}(x):x\in X^{\mathrm{in}},k\in\mathbb{N}\}\subseteq\operatorname{dom}(g)$ ensure the existence of $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $g$. Therefore, we introduce the concept of $(X^{\mathrm{in}},T,\varphi)$-compatibility.
Definition 11 ($(X^{\mathrm{in}},T,\varphi)$-compatibility).
A function $g\in\mathcal{PSD}$ such that $\{T^{k}(x):x\in X^{\mathrm{in}},k\in\mathbb{N}\}\subseteq\operatorname{dom}(g)$ is said to be $(X^{\mathrm{in}},T,\varphi)$-compatible at $k\in{\mathrm{R}}_{\nu}^{s}$ if and only if there exist $\varepsilon>0$ and $\eta>0$ such that for all $x\in X^{\mathrm{in}}$ and for all $j\in\mathbb{N}$, verifying $\varphi(T^{j}(x))>\nu_{k}-\eta$, we have $g(T^{j}(x))>\varepsilon$.
We will say that $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible if there exists some $k\in{\mathrm{R}}_{\nu}^{s}$ at which $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible.
To be $(X^{\mathrm{in}},T,\varphi)$-compatible at $k\in{\mathrm{R}}_{\nu}^{s}$ implies that if $T^{k}(x)$ is $\eta$-optimal for $\varphi$ then $g(T^{k}(x))$ is strictly positive. We extend the strict positivity of $g$ to all orbits of $T$ starting in $X^{\mathrm{in}}$ greater than $\nu_{k}-\eta$.
Proposition 23.
If $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible at $k\in{\mathrm{R}}_{\nu}^{s}$, it is $(X^{\mathrm{in}},T,\varphi)$-compatible at every $j\in{\mathrm{R}}_{\nu}^{s}$ such that $\nu_{j}\geq\nu_{k}$.
Proof.
Suppose that $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible at $k\in\mathbb{N}$. Let $\varepsilon>0$ and $\eta>0$ such that for all $x\in X^{\mathrm{in}}$ and for all $j\in\mathbb{N}$, verifying $\varphi(T^{j}(x))>\nu_{k}-\eta$, we have $g(T^{j}(x))>\varepsilon$. Let $j\in\mathbb{N}$ such that $\nu_{j}\geq\nu_{k}$. Let $x\in X^{\mathrm{in}}$, and $l\in\mathbb{N}$ such that $\varphi(T^{l}(x))>\nu_{j}-\eta$. This implies that $\varphi(T^{l}(x))>\nu_{j}-\eta\geq\nu_{k}-\eta$ and thus $g(T^{l}(x))>\varepsilon$. ∎
The existence of $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $g$ is actually equivalent to the $(X^{\mathrm{in}},T,\varphi)$-compatibility of $g$.
Theorem 8.
A function $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible if and only if there exists $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $g$.
Proof.
Suppose that there exists $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ such that $\alpha(\nu_{k})>0$ for some $k\in{\mathrm{R}}_{\nu}^{s}$ and $\alpha(T^{k}(x))\leq g(T^{k}(x))$ for all $k\in\mathbb{N}$ all $x\in X^{\mathrm{in}}$. Let $k\in{\mathrm{R}}_{\nu}^{s}$ such that $\alpha(\nu_{k})>0$. From Prop. 23, we can suppose that $\nu_{k}>\inf_{n\in\mathbb{N}}\nu_{n}$ and so there exists $\delta>0$ such that $(\nu_{k}-\delta,\nu_{k})\subset I_{X^{\mathrm{in}},T}^{\varphi}$. As $\alpha$ is continuous and strictly increasing on $I_{X^{\mathrm{in}},T}^{\varphi}$, for $\varepsilon:=\alpha(\nu_{k})/2$, there exists $\eta\in(0,\delta]$, such that $0<\nu_{k}-s\leq\eta$ implies that $\varepsilon<\alpha(s)<\alpha(\nu_{k})+\varepsilon$. Now let us consider $j\in\mathbb{N}$ and $x\in X^{\mathrm{in}}$ such that $\varphi(T^{j}(x))>\nu_{k}-\eta$. We have either $\nu_{k}>\varphi(T^{j}(x))$ or $\varphi(T^{j}(x))\geq\nu_{k}$. In the first case, we have, by hypothesis, $g(T^{j}(x))\geq\alpha(\varphi(T^{j}(x)))>\varepsilon$. In the second case, as $\alpha$ is increasing and $g(T^{j}(x))\geq\alpha(\varphi(T^{j}(x)))\geq\alpha(\nu_{k})>\varepsilon$.
Now, assume that $g$ is $(X^{\mathrm{in}},T,\varphi)$-compatible at some $k\in{\mathrm{R}}_{\nu}^{s}$. Then, from Prop.23, $g$ is also $(X^{\mathrm{in}},T,\varphi)$-compatible at $k=\mathbf{K}_{\nu}$. Let $\varepsilon>0$ and $\eta>0$ such that for all $x\in X^{\mathrm{in}}$ and for all $j\in\mathbb{N}$, verifying $\varphi(T^{j}(x))>\nu_{\rm opt}-\eta$, we have $g(T^{j}(x))>\varepsilon$. Then we can define : $\alpha:s\mapsto s-\nu_{\rm opt}+\min\{\eta,\varepsilon\}$. Let $x\in X^{\mathrm{in}}$ and $j\in\mathbb{N}$. Then either $\varphi(T^{j}(x))\leq\nu_{\rm opt}-\min\{\eta,\varepsilon\}$ or $\varphi(T^{j}(x))>\nu_{\rm opt}-\min\{\eta,\varepsilon\}$. For the first case, $\alpha(\varphi(T^{j}(x)))\leq 0\leq g(T^{j}(x)))$. For the second, $0<\alpha(\varphi(T^{j}(x)))\leq\min\{\eta,\varepsilon\}$. By assumption on $g$, $g(T^{j}(x))>\varepsilon$. Note that the set $\{(x,j)\in X^{\mathrm{in}}\times\mathbb{N}:\varphi(T^{j}(x))>\nu_{\rm opt}-\min\{\eta,\varepsilon\}\}$ is nonempty. Finally, as $\alpha$ is strictly increasing, for all $k\in{\mathrm{R}}_{\nu}^{s}$ such that $\nu_{k}\geq\nu_{\rm opt}-\min\{\eta,\varepsilon\}$, we have $\alpha(\nu_{k})>0$. Note that the set $\{k\in\mathbb{N}:\nu_{k}\geq\nu_{\rm opt}-\min\{\eta,\varepsilon\}\}$ is nonempty as $[\mathbf{K}_{\nu},\mathbf{K}^{s}_{\nu}]\cap\mathbb{N}$ is included in it. ∎
Example 9.
Let consider the Opt-Lyapunov function $V$ of Example 5. We prove that $V$ is $(X^{\mathrm{in}},T,\varphi)$-compatible.
We define for all $s\in\mathbb{R}$, $\alpha(s):=s-p^{2}/3+\min\{\zeta,\varepsilon\}$ where $0<\zeta<p^{2}/3-\max\{\nu_{k}:k\notin\operatorname{Argmax}(\nu)\}$ and $\varepsilon:=\min\{\varepsilon_{k},k\in\operatorname{Argmax}{\nu}\}$ where $\varepsilon_{k}=\inf\{V(T^{k}(x)):x\in X^{\mathrm{in}}\}$. The scalar $\varepsilon$ is strictly positive as the function $[\mu,1/2]\ni y\mapsto V(T^{k}((2,1)y)$ is continuous and strictly positive for all $k\in\mathbb{N}$ and $\operatorname{Argmax}(\nu)$ is a finite set of integers. The function $\alpha$ is clearly strictly increasing and continuous on $\mathbb{R}$.
We have for all $j\in\mathbb{N}$ and for all $x\in X^{\mathrm{in}}$, $\alpha(\varphi(T^{j}(x)))=\varphi(T^{j}(x))-p^{2}/3+\min\{\zeta,\varepsilon\}<\varphi(T^{j}(x))-p^{2}/3+p^{2}/3-\max\{\nu_{k}:k\notin\operatorname{Argmax}(\nu)\}$. Then $\alpha(\varphi(T^{j}(x)))$ is strictly negative for all $j\notin\operatorname{Argmax}(\nu)$. Now let $j\in\operatorname{Argmax}(\nu)$. Suppose that $\varphi(T^{j}(x))>p^{2}/3-\min\{\zeta,\varepsilon\}$. As we have $\varphi(T^{j}(x))\leq p^{2}/3$, we get $\alpha(\varphi(T^{j}(x)))\leq\varepsilon\leq V(T^{k}(x))$ for all $k\in\operatorname{Argmax}(\nu)$ and all $x\in X^{\mathrm{in}}$. This leads to $\alpha(\varphi(T^{j}(x)))\leq V(T^{j}(x))$. On the other hand, $\alpha(\varphi(T^{j}(x)))\leq 0$ for all $j\in\operatorname{Argmax}(\nu)$ and $x\in X^{\mathrm{in}}$ such that $\varphi(T^{k}(x))\leq p^{2}/3-\min\{\eta,\varepsilon\}$. We have proved that $\alpha(\varphi(T^{k}(x)))\leq V(T^{k}(x))$ for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$. Moreover, as the set of $\{x\in X^{\mathrm{in}}:\varphi(T^{j}(x))>p^{2}/3-\min\{\eta,\varepsilon\}\}$ is nonempty for all $j\in\operatorname{Argmax}(\nu)$, we conclude that for all $j\in\operatorname{Argmax}(\nu)$, there exists $x\in X^{\mathrm{in}}$ such that $0<\alpha(\varphi(T^{j}(x)))$ and as $\alpha$ is increasing $0<\alpha(\nu_{j})$ for all $j\in\operatorname{Argmax}(\nu)$ and $\alpha$ is a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $V$ making $V$ $(X^{\mathrm{in}},T,\varphi)$-compatible.
Actually, by simple computations, we have $\varepsilon=8\mu^{2}(2/3)^{2\overline{n}-1}$. One can compute $\max\{\nu_{k}:k\notin\operatorname{Argmax}(\nu)\}$ as follows. Let $n\in\mathbb{N}$. We have, for all $y\in\mathbb{R}_{+}^{*}$, $\varphi_{n+1}(y)-\varphi_{n}(y)=(3/2)^{n}y(-(15/4)(3/2)^{n}y+p)$. Thus $\varphi_{n+1}(y)-\varphi_{n}(y)=(3/2)^{n}y(-(15/4)(3/2)^{n}y+p)$ is strictly negative if and only if $n\geq n_{y}^{*}:=\max\{\lfloor\ln(4p/(15y))/\ln(3/2)\rfloor+1,0\}$. Then for all $y\in\mathbb{R}_{+}^{*}$, $\mathbb{N}\ni n\mapsto\varphi_{n}(y)$ is maximized at $n_{y}^{*}$. We deduce that $n_{1/2}^{*}=\lfloor\ln(8p/15)/\ln(3/2)\rfloor+1$ and, as $\mu\in(0,1/3)$, $n_{\mu}^{*}=\lfloor\ln(4p/(15\mu))/\ln(3/2)\rfloor+1$. This implies that $n_{1/2}^{*}\leq\underline{n}-1$ and thus we have $\nu_{n}\leq\nu_{n_{1/2}^{*}}$ for all $n\leq\underline{n}-1$. This also implies that $n_{\mu}^{*}<\overline{n}$ and so $\nu_{n}<\nu_{\overline{n}}$ for all $n>\overline{n}$. Finally, $\max\{\nu_{k}:k\notin\operatorname{Argmax}(\nu)\}=\max\{\varphi_{n_{1/2}^{*}}(1/2),\varphi_{\overline{n}}(\mu)\}$.
5.2.4 Direct and Converse Opt-Lyapunov Theorems
We present both direct and converse Opt-Lyapunov theorems using the $(X^{\mathrm{in}},T,\varphi)$-compatibility notion. More precisely, the direct version aims to prove that if a $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov function exists (and such that $\overline{N_{T}}(V)>0$), then we can construct a couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for $\nu$. The converse version exhibits a $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov function when a couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ is useful for $\nu$. Moreover, we are able to produce a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for this Opt-Lyapunov function.
We remark that from Prop. 14, if $V$ is an Opt-Lyapunov for $(X^{\mathrm{in}},T)$ then $\{T^{k}(x):k\in\mathbb{N},x\in X^{\mathrm{in}}\}\subseteq\operatorname{dom}(V)$.
Proposition 24.
Assume that there exists a $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov $V$ such that $N_{V}(T)=0$. Then $\mathbf{K}^{s}_{\nu}=0$ and $\nu_{\rm opt}=\overline{\varphi_{X^{\mathrm{in}}}}$.
Proof.
If $N_{V}(T)=0$ then for all $x\in\mathbb{R}^{d}$ such that $V(x)>0$, $V(T(x))=0$. This also implies that $V(T^{k}(x))=0$ for all $k\in\mathbb{N}^{*}$. Let $\alpha$ be a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility. Thus, for all $k\in\mathbb{N}^{*}$, $\alpha(\varphi(T^{k}(x)))\leq V(T^{k}(x))=0$ and as $\alpha$ is increasing and continuous, $\alpha(\nu_{k})\leq 0$ for all $k\in\mathbb{N}^{*}$. As for some $j\in{\mathrm{R}}_{\nu}^{s}$, $\alpha(\nu_{j})>0$, this leads to $j=0$ and as $\alpha$ is strictly increasing and continuous on $I_{X^{\mathrm{in}},T}^{\varphi}$ then for all $k\in\mathbb{N}^{*}$, $\nu_{0}>\nu_{k}$ and $\mathbf{K}^{s}_{\nu}=0$. This also implies that $\nu_{\rm opt}=\overline{\varphi_{X^{\mathrm{in}}}}$. ∎
Theorem 9 (Direct Opt-Lyapunov Theorem).
Assume that there exists a $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov $V$ such that $N_{V}(T)>0$. Let $\alpha$ be a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $V$. Let $J=\overline{\operatorname{\operatorname{conv}}}(I_{X^{\mathrm{in}},T}^{\varphi}\cup I)$ where $\alpha(I)=[0,1]$. We define the function $h:x\mapsto\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(x\overline{V_{X^{\mathrm{in}}}}\right)$. Then $(h,N_{V}(T))\in\mathbb{\Gamma}(\nu)$ is useful for $\nu$.
Proof.
The function ${{\alpha}_{\left(J\right)}^{-1}}$ is strictly increasing and continuous. Thus, we have, for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$, $\varphi(T^{k}(x))\leq\displaystyle{{\alpha}_{\left(J\right)}^{-1}}(V(T^{k}(x))$. By Def. 9 and Prop. 17, $V(T^{k}(x))\leq(N_{V}(T))^{k}V(x)\leq N_{V}(T)^{k}\overline{V_{X^{\mathrm{in}}}}$. As ${{\alpha}_{\left(J\right)}^{-1}}$ is strictly increasing, we conclude that $\nu_{k}=\sup_{x\in X^{\mathrm{in}}}\varphi(T^{k}(x))\leq\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(N_{V}(T)^{k}\overline{V_{X^{\mathrm{in}}}}\right)$. Note that $x\mapsto\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(x\overline{V_{X^{\mathrm{in}}}}\right)$ is strictly increasing and continuous on $J\supset[0,1]$ as ${{\alpha}_{\left(J\right)}^{-1}}$ is and $\overline{V_{X^{\mathrm{in}}}}$ is not zero. Moreover, by Prop. 18 and hypothesis, $N_{V}(T)\in(0,1)$. As $\alpha$ is a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $V$, there exists $j\in{\mathrm{R}}_{\nu}^{s}$ such that $\alpha(\nu_{j})>0$. As ${{\alpha}_{\left(J\right)}^{-1}}$ is strictly increasing and $\alpha(I)\ni 0$, $\nu_{j}>\displaystyle{{\alpha}_{\left(J\right)}^{-1}}(0)=\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\left(0\times\overline{V_{X^{\mathrm{in}}}}\right)=h(0)$. ∎
Example 10.
Following Th. 9, from the Opt-Lyapunov function $V$ defined at Example 5 and the ($X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $V$, $\alpha$ defined at Example 9, the function $h:[0,1]\ni s\mapsto\alpha_{\mathbb{R}}^{-1}(s\overline{V_{X^{\mathrm{in}}}})$ belongs to $\mathbb{\Gamma}_{\rm fun}(\nu)$ and $\overline{N_{T}}(V)=4/9\in\mathbb{\Gamma}_{\rm sc}(\nu,h)$. Then $h_{V}(s)=s+p^{2}/3-\min\{\zeta,\varepsilon\}$ with $\varepsilon=8\mu^{2}(2/3)^{2\overline{n}-1}$ and $\zeta\in(0,p^{2}/3-\max\{\varphi_{n_{1/2}^{*}}(1/2),\varphi_{\overline{n}}(\mu)\})$ where $n_{1/2}^{*}$ is the maximizer for the function $n\mapsto\varphi_{n}(1/2)$.
We draw a table as we done at Example 3 to compare the direct (from the function $h$ of Example 3) and the Opt-Lyapunov approaches. We have some flexibility about $\zeta$ so we propose two values for a comparison purpose. Hence, we compare the computations for $h_{\zeta_{1}}$ associated to $\zeta_{1}=(p^{2}/3-\max\{\varphi_{n_{1/2}^{*}}(1/2),\varphi_{\overline{n}}(\mu)\})/2$ and $h_{\zeta_{2}}$ associated to $\zeta_{2}=(p^{2}/3-\max\{\varphi_{n_{1/2}^{*}}(1/2),\varphi_{\overline{n}}(\mu)\})/1000$. The evaluations for $\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{1}},4/9)\rfloor$ and for $\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{2}},4/9)\rfloor$ are given at Table 3.
$p=30$ $p=9$ $p=3$
$\mu$
$\frac{1}{3}$
$\mu$
$\frac{1}{10}$
$\mu$
$\frac{1}{1000}$
$\mu$
$\frac{1}{3}$
$\mu$
$\frac{1}{10}$
$\mu$
$\frac{1}{1000}$
$\mu$
$\frac{1}{3}$
$\mu$
$\frac{1}{10}$
$\mu$
$\frac{1}{1000}$
$\mathbf{K}^{s}_{\nu}$ 8 11 22 5 8 19 2 3 17
$\lfloor\mathfrak{F}_{\nu}(k,a,\beta)\rfloor$ 9 12 23 7 9 20 3 7 18
$\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{1}},\frac{4}{9})\rfloor$ 8 14 36 5 11 33 4 8 31
$\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{2}},\frac{4}{9})\rfloor$ 8 14 36 9 11 33 12 12 31
Table 3: Evaluations of the function $\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{i}},\beta)\rfloor$ described at Eq. (4) for the sequence $\nu$ of Eq. (2) and two pairs $(h_{\zeta_{1}},4/9)$ and $(h_{\zeta_{2}},4/9)$ of Example 10.
The evaluations of $\lfloor\mathfrak{F}_{\nu}(k,h_{\zeta_{i}},4/9)\rfloor$ from the Opt-Lyapunov function proposed at Example 5 and associated $(X^{\mathrm{in}},T,\varphi)$-certificates of compatibility provide worst upper bounds when $\mu$ is small (equal to $1/10$ or $1/1000$) but the optimal value for $\mu=1/3$. A small value for $\zeta$ ($\zeta_{2}=(p^{2}/3-\max\{\varphi_{n_{1/2}^{*}}(1/2),\varphi_{\overline{n}}(\mu)\})/1000$) makes decrease the precision for $\mu=1/3$. We recall that the y-intercept of the affine function $h_{\zeta}$ is $p^{2}/3-\min\{\varepsilon,\eta\}$ where $\varepsilon$ is a very small value when $\mu$ is small. Hence, when $\mu=1/10$ or $\mu=1/1000$, $\zeta_{1}$ or $\zeta_{2}$ are still greater than $\varepsilon$.
Note that the values $h_{\zeta_{i}}^{-1}(\nu_{k})$ are strictly positive if and only if for $k\in\operatorname{Argmax}(\nu)$. This is sufficient to compute an upper bound of $\mathbf{K}^{s}_{\nu}$. ∎
We establish a ”converse Lyapunov theorem”. The construction of the Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ is inspired by the one proposed by Yoshizawa [Yos66].
Theorem 10 (Converse Opt-Lyapunov Theorem).
Let $(h,\beta)\in\mathbb{\Gamma}(\nu)$ be useful for $\nu$. Let us define $s\mapsto\omega(s):=\min\left\{h(1),\max\left\{s,h(0)\right\}\right\}$. Then, the function $V\in\mathcal{PSD}$ defined as follows:
$x\mapsto V(x)=\sup_{k\in\mathbb{N}}\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega\left(\varphi(T^{k}(x))\right)\right)$
is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$. Moreover, the function $\hat{h}$ defined as follows:
$\hat{h}:\mathbb{R}\mapsto\left\{\begin{array}[]{lr}s-h(0)&\text{ if }s\leq h(0)\\ \displaystyle{{h}_{\left([0,1]\right)}^{-1}}(s)&\text{ if }s\in(h(0),h(1)]\end{array}\right.$
is a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility for $V$ making the function $V$ $(X^{\mathrm{in}},T,\varphi)$-compatible.
Proof.
First, we prove that the function $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
Let us prove that $V$ takes its values in $[0,+\infty]$. Let $x\in\mathbb{R}^{d}$. As we have $h(1)\geq\omega(s)\geq h(0)$, we have, as ${{h}_{\left([0,1]\right)}^{-1}}$ is increasing, $1\geq\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(\omega(\varphi(T^{k}(x))))\geq 0$ for all $k\in\mathbb{N}$. We conclude, as $\beta\in(0,1)$ that $V(x)\geq 0$ and $V(x)=+\infty$ for all $x\in\mathbb{R}^{d}$ such that $\varphi(T^{k}(x))\geq h(1)$ for all $k$ sufficiently large.
By definition of $\mathbb{\Gamma}(\nu)$, we have for all $x\in X^{\mathrm{in}}$ and all $n\in\mathbb{N}$, $\varphi(T^{n}(x))\leq h(\beta^{n})<h(1)$ and so $V(x)\leq\sup_{k\in\mathbb{N}}\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(h(\beta^{k}))=\sup_{k\in\mathbb{N}}\beta^{-k}\beta^{k}=1$. This implies that $\overline{V_{X^{\mathrm{in}}}}\in[0,1]$. By assumption, $(h,\beta)\in\mathbb{\Gamma}(\nu)$ and $\mathcal{S}(\nu,h)$ is nonempty. Let us take $k\in\mathbb{N}$ such that $h(1)>\nu_{k}>h(0)$. So there exists $\varepsilon>0$ such that $\nu_{k}-\varepsilon>h(0)$. For this $\varepsilon$, there exists $x\in X^{\mathrm{in}}$ such that $\nu_{k}\geq\varphi(T^{k}(x))\geq\nu_{k}-\varepsilon$. Finally, for this $x$, $V(x)\geq\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(\varphi(T^{k}(x))>\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(h(0))=0$ as ${{h}_{\left([0,1]\right)}^{-1}}$ strictly increases. We conclude that $\overline{V_{X^{\mathrm{in}}}}\in(0,1]$. Finally, we have to prove that $V\circ T\leq\beta V$. Let $y\in\mathbb{R}^{d}$, if $V(y)=+\infty$, the result holds. Now suppose that $y\in\operatorname{dom}V$, we have:
$\begin{array}[]{rl}V(T(y))=&\displaystyle{\sup_{k\in\mathbb{N}}\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega(\varphi(T^{k+1}(y))\right)}\\ =&\displaystyle{\sup_{j\in\mathbb{N}^{*}}\beta^{-j+1}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega(\varphi(T^{j}(y))\right)}\\ =&\displaystyle{\beta\sup_{j\in\mathbb{N}^{*}}\beta^{-j}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega(\varphi(T^{j}(y))\right)}\\ \leq&\displaystyle{\beta\sup_{j\in\mathbb{N}}\beta^{-j}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega(\varphi(T^{j}(y))\right)=\beta V(y)}\end{array}$
We conclude that $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$.
[…]
6 A Technical Summary
We have proposed a function $\mathfrak{F}_{\nu}$ at Eq. (4) whose the integer part of values provide solutions for Problem 2. This function is based on couples $(h,\beta)\in\mathbb{\Gamma}(\nu)$ useful for $\nu$. Without this condition, $(h,\beta)$ is useless as $\mathfrak{F}_{\nu}$ is equal to +$\infty$+ ∞. When the supremum of $(\nu_{k})_{k}$ is strictly greater than its limit superior, a useful couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ exists (Th. 1) and $h$ can be chosen as an affine function. Furthermore, we can compute the greatest maximizer of $(\nu_{k})_{k}$ from this couple (Th. 3) and thus we can solve Problem 1. Unfortunately, to directly compute a pair $(h,\beta)\in\mathbb{\Gamma}(\nu)$ is as difficult as solving Problem 1 and we have to construct a couple $(h,\beta)\in\mathbb{\Gamma}(\nu)$ from alternative methods. We propose alternatives coming from classical tools from stability theory of difference equations.
The inequality $\varphi(T^{k}(x))\leq h(\beta^{k})$ for all $x\in X^{\mathrm{in}}$ and for all $k\in\mathbb{N}$ can be viewed as a kind of $\mathcal{KL}$ inequality that appears in asymptotic stability studies. Indeed, $h$ is strictly increasing and continuous on $[0,1]$ and $\beta\in(0,1)$ making the sequence $(h(\beta^{k}))_{k}$ strictly decreasing and convergent to $h(0)$. From this observation, we propose a more general $\mathcal{K}\mathcal{L}$ class of functions called $\mathcal{KL}_{\rm gen}$ where the positivity with respect to the state variable and the continuity assumptions are removed (compared with the $\mathcal{K}\mathcal{L}$ class). We have proved (Th. 5) that a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ exists if and only if $\nu_{\rm opt}$ is finite. Then, we extend Sontag’s result to the $\mathcal{KL}_{\rm gen}$ class (Lemma 2 and Lemma 3) to construct $h\in\Omega([0,1])$, useful for $\nu$, from a $\mathcal{KL}_{\rm gen}$ function and such that $(h,e^{-1})\in\mathbb{\Gamma}(\nu)$ (Th. 6). One of the advantage of the Sontag’s result extension is that the scalar part of elements of $\mathbb{\Gamma}(\nu)$ can be fixed and it remains to construct $h$ from the $\mathcal{KL}_{\rm gen}$ certificate. In practice, to decide whether a function $(\gamma,\operatorname{Fin}\left(X^{\mathrm{in}}\right))$ is a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound of $\varphi(T^{k}(x))$ seems difficult to make.
One constructive alternative of $\mathcal{K}\mathcal{L}$ certificates consists in Lyapunov functions. Classical Lyapunov functions satisfy pure functional inequations and are not defined from the orbits of the dynamical system. One of the most interesting theoretical result for Lyapunov functions is the necessity of their existence for the stability (namely converse Lyapunov theorems). For our specific problem, we propose a new type of Lyapunov function called Opt-Lyapunov functions for which we obtain both direct (Th. 9) and converse theorem (Th 10). The main difference between classical and Opt-Lyapunov functions is that Opt-Lyapunov functions are nonnegative but not definite (many zeros) and can take infinite values (outside reachable values taken by the dynamical system). The fact that Opt-Lyapunov functions are not definite (and not continuous) makes classical comparison functions impossible to use. In the classical case, comparison functions serve as link between our objective function $\varphi$ and the classical Lyapunov function. This link (and comparison functions) is replaced by the $(X^{\mathrm{in}},T,\varphi)$-compatibility notion (Th. 8). Theorem 11 sums up the equivalences proved in this paper relatively to the peaks computation problem.
Theorem 11 (Summary).
The following statements are equivalent:
  1. 1.
    The set ${\mathrm{R}}_{\nu}^{s}$ is nonempty;
  2. 2.
    There $(h,\beta)\in\mathbb{\Gamma}(\nu)$ for which $\mathcal{S}(\nu,h)$ is nonempty;
  3. 3.
    There exists $(h,\beta)\in\mathbb{\Gamma}(\nu)$ such that $\mathfrak{F}_{\nu}(\mathbf{K}^{s}_{\nu},h,\beta)$ is finite and $\nu_{\rm opt}=\max_{k\in\mathbb{N}}\{\nu_{k}:k=0,\ldots,\mathbf{K}^{s}_{\nu}\}$;
  4. 4.
    There exists $\theta\in\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ and $\gamma\in\mathcal{KL}_{\rm gen}$ such that $\varphi(T^{k}(x))\leq\gamma(\theta(x),k)$ for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$ and $\inf_{t\geq 0}\gamma(\overline{\theta_{X^{\mathrm{in}}}},t)<\nu_{n}$ for some $n\in\mathbb{N}$;
  5. 5.
    There exists an Opt-Lyapunov $(X^{\mathrm{in}},T,\varphi)$ compatible function.
The proof of Theorem is done by sketching the functions constructed in the proofs of earlier theorems. Figure 1 provides a graphical view of different results involved to prove Th. 11. In Figure 1, the notation $\mathcal{OL}_{\rm comp}$ stands for the set of $(X^{\mathrm{in}},T,\varphi)$-compatible Opt-Lyapunov functions.
[…]
Figure 1: The scheme showing the mappings between $\mathcal{KL}_{\rm gen}$, $\mathcal{OL}_{\rm comp}$ and $\mathbb{\Gamma}(\nu)$
As $V\in\mathcal{OL}_{\rm comp}$ is $(X^{\mathrm{in}},T,\varphi)$-compatible, we associate with them a $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility. For a such a function $V$, we denote by $\alpha_{V}$ the chosen $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility.
We now clarify the circled numbers that appear in Figure 1.
[…]
    $\begin{array}[]{ccc}\mathbb{\Gamma}(\nu)&\longrightarrow&\mathcal{KL}_{\rm gen}\times\operatorname{Fin}\left(X^{\mathrm{in}}\right)\\ (h,\beta)&\longmapsto&\left(\mathbb{R}\times\mathbb{R}_{+}\ni(s,t)\mapsto\displaystyle{\min\{s,h(\beta^{t})\}},\displaystyle{\mathbb{R}^{d}\ni x\mapsto\sup_{k\in\mathbb{N}}\varphi(T^{k}(x))}\right)\end{array}$
[…]
    $\begin{array}[]{ccc}\mathcal{KL}_{\rm gen}\times\operatorname{Fin}\left(X^{\mathrm{in}}\right)&\longrightarrow&\mathbb{\Gamma}(\nu)\\ (\gamma,\theta)&\longmapsto&(h,e^{-1})\end{array}$
    The function $h$ cannot be explicitly formulated and is constructed following the proofs of Lemma 2 and Lemma 3.
[…]
    $\begin{array}[]{ccccc}\mathcal{OL}_{\rm comp}&\longrightarrow&\mathcal{OL}_{\rm comp}\times\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}&\longrightarrow&\mathbb{\Gamma}(\nu)\\ V&\longmapsto&(V,\alpha_{V})&\longmapsto&(h_{V},\overline{N_{T}}(V))\end{array}$
    where $h_{V}$ is defined as follows:
    $[0,1]\ni s\mapsto h_{V}(s)=\displaystyle{{\alpha_{V}}_{\left(J\right)}^{-1}}(s\overline{V_{X^{\mathrm{in}}}})$
    where $J=\overline{\operatorname{\operatorname{conv}}}(I_{X^{\mathrm{in}},T}^{\varphi}\cup I)\in\mathrm{SIC}(\alpha_{V})$ where $\alpha_{V}(I)=[0,1]$.
[…]
    $\begin{array}[]{ccc}\mathbb{\Gamma}(\nu)&\longrightarrow&\mathcal{OL}_{\rm comp}\times\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}\\ (h,\beta)&\longmapsto&(V,\alpha_{V})\end{array}$
    where:
    • $\bullet$
      ${\mathbb{R}^{d}\ni x\mapsto V(x)=\sup_{k\in\mathbb{N}}\beta^{-k}\displaystyle{{h}_{\left([0,1]\right)}^{-1}}\left(\omega(\varphi(T^{k}(x)))\right)}$;
    • $\bullet$
      ${\mathbb{R}\ni s\mapsto\omega(s)=\min\{h(1),\max\{s,h(0)\}\}}$;
    • $\bullet$
      $(-\infty,h(1)]\ni s\mapsto\alpha_{V}(s)=\left\{\begin{array}[]{lr}s-h(0)&\text{ if }s\leq h(0)\\ \displaystyle{{h}_{\left([0,1]\right)}^{-1}}(s)&\text{ if }s\in[h(0),h(1)]\end{array}\right.$
Appendix
Proof of Lemma 4.
Lemma 4.
Let $\gamma$ be strictly greater than 1. Let $f\in\mathcal{K}_{\infty}$ verifying for all $x>0$, $f(x)>x$. Then, there exists a function $g\in\mathcal{K}$ which satisfies $g(f(x))=\gamma g(x)$ for all $x\in\mathbb{R}_{+}$.
Proof.
As we have $f(x)>x$ for all $x>0$, we have $f(1)>1$ and thus the interval $[1,f(1))$ is nonempty. We define $h:[1,f(1))$ continuous, strictly increasing whose values are strictly positive such that $h(1)=1$ and $\lim_{x\mapsto f(1)^{-}}h(x)=\gamma$. We could take for example $h$ the affine map linking $(1,1)$ to $(f(1),\gamma)$. Now, as the sequence $(f^{k}(1))_{k}$ is strictly increasing and unbounded (if $M=\sup_{k\in\mathbb{N}}f^{k}(1)>0$ is finite then as $f$ is increasing and continuous $f(M)=M$ but $M<f(M)$ by assumption on $f$). As $f$ is invertible on $\mathbb{R}_{+}$ and we have $f^{-1}(x)<x$ for all $x>0$, the sequence $(f^{-k}(1))_{k}$ is strictly decreasing and positive and converges to 0 as $k$ tends to +$\infty$+ ∞. Hence $\mathbb{R}_{+}^{*}$ is union of disjoint intervals $[f^{k}(1),f^{k+1}(1))$ where $k\in\mathbb{Z}$. Then for all $x>0$, there exists a unique $k\in\mathbb{Z}$ such that $x\in[f^{k}(1),f^{k+1}(1))$. We define a function $g$ on $\mathbb{R}_{+}$ as follows $g(0)=0$ and for all $x>0$, $g(x):=\gamma^{k}h(f^{-k}(x))$ where $k$ is the unique integer such that $x\in[f^{k}(1),f^{k+1}(1))$. As $g$ is equal to $h$ on $[1,f(1)]$, then $g$ is strictly increasing, continuous and strictly positive on $[1,f(1)]$. Then, on the open interval $(f^{k}(1),f^{k+1}(1))$ for $k\neq 0$, $g$ is the composition $t\mapsto\gamma^{k}t$, $h$ and $f^{-k}$ and thus is strictly increasing, continuous and strictly positive on $(f^{k}(1),f^{k+1}(1))$ for all $k\in\mathbb{Z}$. Now, for all $k\neq 0$, $\lim_{y\to{f^{k}(1)}^{-}}g(y)=\lim_{y\to{f^{k}(1)}^{-}}\gamma^{k-1}h(f^{-k+1}(y))=\lim_{z=f^{-k}(y)\to 1^{-}}\gamma^{k-1}h(f(z))=\gamma^{k-1+1}=\gamma^{k}$. We also have $\lim_{y\to{f^{k}(1)}^{+}}g(y)=\lim_{y\to{f^{k}(1)}^{+}}\gamma^{k}h(f^{-k}(y))=\lim_{z\to 1^{-}}\gamma^{k}h(z)=\gamma^{k}h(1)=\gamma^{k}$. We conclude that $g$ is continuous on $\mathbb{R}_{+}^{*}$. Now, we prove that $g$ is continuous at 0 (at right). Let us consider a sequence $(x_{n})_{n\in\mathbb{N}}$ of positive reals which converges to 0. Let $n\in\mathbb{N}$. There exists a unique $k_{n}\in\mathbb{Z}$ such that $f^{k_{n}}(1)\leq x_{n}\leq f^{k_{n}+1}(1)$. We get $1\leq f^{-k_{n}}(x_{n})\leq f(1)$ and as $h$ is increasing on $[1,f(1))$ and continuous (at left) at $f(1)$, $1=h(1)\leq h(f^{-k_{n}}(x_{n}))\leq h(f(1))=\gamma$. Finally, $\gamma^{k_{n}}\leq g(x_{n})=\gamma^{k_{n}}h(f^{-k_{n}}(x_{n}))\leq\gamma^{k_{n}+1}$. As $(x_{n})_{n\in\mathbb{N}}$ tends to 0, the sequence $(k_{n})_{n\in\mathbb{N}}$ tends to $\infty$- ∞ and we conclude that $(g(x_{n}))_{n\in\mathbb{N}}$ tends to $0=g(0)$ and $g$ is continuous at 0. We have already proved that $g$ is strictly increasing on each interval $[f^{k}(1),f^{k+1}(1))$. It is easy to see that it is also the case when we pick two reals $x,y$ in two different intervals : $x<y$ implies $g(x)<g(y)$.
Finally, we have to prove that $g(f(x))=\gamma g(x)$ for all $x\in\mathbb{R}_{+}$. As $f(0)=0$, this holds for $x=0$. Now let us take $x>0$. This real $x$ belongs to some interval $[f^{k}(1),f^{k+1}(1))$. Thus, $f(x)\in[f^{k+1}(1),f^{k+2}(1))$ and $g(f(x))=\gamma^{k+1}h(f^{-k-1}(f(x))=\gamma\gamma^{k}h(f^{-k}(x))=\gamma g(x)$. This ends the proof. ∎
Proof of Lemma 5
Lemma 5.
Let $\lambda\in(0,1)$. Let $f\in\mathcal{K}$ verifying for all $x>0$, $f(x)<x$. Then, there exists a function $g\in\mathcal{K}$ satisfying $g(f(x))=\lambda g(x)$ for all $x\in\mathbb{R}_{+}$.
Proof.
The function $f^{-1}$ is continuous, strictly increasing and definite and satisfies $f^{-1}(x)>x$ for all $x>0$. Then using Lemma 4, there exists $g\in\mathcal{K}$ such that $g(f^{-1}(x))=\lambda^{-1}g(x)$ for all $x\in\mathbb{R}$. Then for all $y\in\mathbb{R}_{+}$, there exists a unique $x\in\mathbb{R}_{+}$ such that $y=f^{-1}(x)$ or equivalently $x=f(y)$. Then, we have $g(y)=\lambda^{-1}g(f(y))$ and we conclude that $g(f(y))=\lambda g(y)$ for all $y\in\mathbb{R}_{+}$. ∎